### v1

# GraphXForm: Graph transformer for computer-aided molecular design with application to extraction

Abstract
Generative deep learning has become pivotal in molecular design for drug discovery and materials science. A widely used paradigm is to pretrain neural networks on string representations of molecules and fine-tune them using reinforcement learning on specific objectives. However, string-based models face challenges in ensuring chemical validity and enforcing structural constraints like the presence of specific substructures. We propose to instead combine graph-based molecular representations, which can naturally ensure chemical validity, with transformer architectures, which are highly expressive and capable of modeling long-range dependencies between atoms. Our approach iteratively modifies a molecular graph by adding atoms and bonds, which ensures chemical validity and facilitates the incorporation of structural constraints. We present GraphXForm, a decoder-only graph transformer architecture, which is pretrained on existing compounds and then fine-tuned using a new training algorithm that combines elements of the deep cross-entropy method with self-improvement learning from language modeling, allowing stable fine-tuning of deep transformers with many layers. We evaluate GraphXForm on two solvent design tasks for liquid-liquid extraction, showing that it outperforms four state-of-the-art molecular design techniques, while it can flexibly enforce structural constraints or initiate the design from existing molecular structures.
Machine Learning, ICML
[…]
1 Introduction
Molecular design plays an important role across many fields, such as drug discovery, materials science, and chemical engineering. The immense chemical search space, estimated to contain between $10^{60}$ and $10^{100}$ potential molecules (Schneider & Fechner, 2005), renders manual approaches to molecular design both arduous and resource-intensive.
Advancements in deep learning have significantly impacted molecular design, enabling efficient navigation of the chemical space with the help of neural networks (Fu et al., 2022; Gómez-Bombarelli et al., 2018; Segler et al., 2018; Bjerrum & Threlfall, 2017; Zhang et al., 2023a). A prevalent paradigm is to represent molecules as strings of text, such as SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2020), and use neural network architectures from language modeling, such as recurrent neural networks (RNNs) or transformers (Vaswani, 2017), to generate novel molecular structures. These architectures are typically pretrained on large datasets of existing molecules to learn general underlying patterns in the strings and then fine-tuned on specific objective functions via reinforcement learning (RL) for downstream tasks (Olivecrona et al., 2017; Gao et al., 2022; Xu et al., 2024; Mazuz et al., 2023; Gupta et al., 2018).
To address the challenge of capturing long-range dependencies in sequential data that RNNs face, transformers (Vaswani, 2017) have been successfully applied due to their ability to model long-range dependencies more efficiently and as they are widely used in language modeling today (Xu et al., 2024). However, transformers are resource-intensive, and needing to fine-tune them with RL can limit the size that is used in practice (Xu et al., 2024; Henderson et al., 2018; Schulman et al., 2017). In general, chemical language models may propose string representations of molecules with invalid chemical structures – for example, when SMILES syntax or valence constraints are violated – which has led to numerous works aimed at circumventing this problem (O’Boyle & Dalke, 2018; Krenn et al., 2020; Cheng et al., 2023; Dai et al., 2018). Despite recent evidence that invalid SMILES can actually be beneficial from a language modeling perspective (Skinnider, 2024), they can harm the RL component of the pipeline as they can increase sample complexity and necessitate reward shaping to account for them. Additionally, the sequential nature of string synthesis makes it difficult to enforce structural constraints, such as ensuring the presence of a minimum number of specific atom types, restrictions on bonds, or even enforcing the presence of certain substructures within the molecules.
An alternative approach is to represent molecules directly as graphs, where atoms are nodes and bonds are edges, and to develop models that modify a molecular graph or design it directly (Wang et al., 2023; Li et al., 2018; Jin et al., 2018; Zhou et al., 2019; Jensen, 2019; De Cao & Kipf, 2018; Zang & Wang, 2020; Zhang et al., 2023b; Mahmood et al., 2021; Maziarka et al., 2019; Rittig et al., 2023d). Working at the graph level, allows explicit encoding of atomic interactions and bonding rules, ensuring chemical validity, and makes it straightforward to start from existing structures and modify them. For example, graph-based methods like GraphGA (Jensen, 2019) employ genetic algorithms to modify molecular graphs directly, even outperforming several neural-based string-synthesis methods without relying on neural networks. As a deep learning example, Zhou et al. (2019) use deep Q-Learning (Mnih, 2013) with a simple feedforward network to optimize graphs from scratch by adding or removing atoms and bonds.
We propose to combine and extend the strengths of both paradigms: employing transformers for their ability to capture long-range dependencies in sequence data, and leveraging pretraining on existing molecules, all while working directly on the molecular graph. More specifically, we use graph transformers (Ying et al., 2021; Maziarka et al., 2019) and formulate molecule design as a sequential task, where a molecular graph – starting from an arbitrary structure – is iteratively extended by placing atoms, bonding them to existing atoms, and modifying the graph by increasing bond orders.
To this end, we introduce GraphXForm, a decoder-only graph transformer architecture that guides these incremental decisions, predicting the next action based on the current molecular graph. Pretrained on existing compounds, GraphXForm employs a novel fine-tuning approach that combines elements of the deep cross-entropy method (Wagner, 2021) and self-improvement learning (Huang et al., 2023; Pirnay & Grimm, 2024a). Unlike commonly used deep reinforcement learning methods like REINFORCE (Williams, 1992), this approach allows for stable training of deep transformers with many layers.
We test GraphXForm for the design of solvents, which are highly relevant for chemical processes. While algorithmic advances in molecular design have primarily focused on de novo drug development, other areas such as materials science and chemical engineering have only recently gained attention. Recent applications include catalyst (Schilter et al., 2023), fuel (Rittig et al., 2023d; Sarathy & Eraqi, 2024), polymer (Jiang et al., 2024; Yue et al., 2024), surfactant (Nnadili et al., 2024), chemical reaction substrate (Nigam et al., 2023), and solvent (König-Mattern et al., 2024) design. We focus specifically on solvents due to their vital role in industrial chemical processes such as reactions, separations, and extractions. While König-Mattern (König-Mattern et al., 2024) recently applied a graph-based genetic algorithm for solvent design, the use of generative ML-based approaches remains underexplored. Therefore, our goal is to compare newer design methods and expand their capabilities.
We evaluate GraphXForm on two liquid-liquid extraction tasks. The objective function in these tasks is defined by a separation factor based on activity coefficients at infinite dilution. We compare GraphXForm on these two downstream tasks to four state-of-the-art molecular design approaches (Graph GA (Jensen, 2019), REINVENT-Transformer (Xu et al., 2024), Junction Tree VAE (Jin et al., 2018), and STONED (Nigam et al., 2021)). Additionally, we demonstrate GraphXForm’s unique flexibility and stability by incorporating structural constraints conceptually suited for solvent design, such as preventing certain bond types or preserving complex molecular substructures, allowing us to design solvents with highly tailored properties. This capability highlights the strength of our approach in tackling design tasks that are difficult or infeasible for existing methods.
Our contributions are summarized as follows:
  • We formulate molecular design as a sequential task, where an initial structure (e.g., a single atom) is iteratively modified by adding atoms and bonds.
  • We introduce a graph transformer architecture that takes a molecular graph as input and outputs probability distributions for atom placement and bond modification. This approach maintains the notion of using transformers for molecular design while moving away from string-based methods.
  • We propose a training algorithm that enables the stable and efficient fine-tuning of deep graph transformers on downstream tasks.
  • We demonstrate that our method outperforms state-of-the-art molecular design techniques on two solvent design tasks and can be easily adapted to satisfy structural constraints or preserve specific molecular structures.
Our code for pretraining and fine-tuning is available at:
https://github.com/grimmlab/graphxform.
2 Methods and Modeling
In this section, we introduce GraphXForm. In Section 2.1, we formally set up the molecular design task as the sequential construction of a graph. As in a deep reinforcement learning setup, the goal is to find a policy network that guides these sequential decisions. In Section 2.2, we introduce our algorithm for training the policy network, before describing the used transformer architecture in Section 2.3.
2.1 Molecular design
Molecular graph
In the following, we represent a molecule by its hydrogen-suppressed graph representation, where nodes correspond to atoms and edges correspond to bonds. For ease of notation, we assume an arbitrary ordering over the nodes. Let $\Sigma=(\Sigma_{1},\dots,\Sigma_{k})$ be an underlying alphabet of possible atom types. We represent a molecule with $n$ atoms by a pair $(\bm{a},\bm{B})$, where $\bm{a}=(a_{1},\dots,a_{n})\in\{1,\dots,k\}^{n}$ and $a_{i}$ indicates that the $i$-th node is of atom type $\Sigma_{a_{i}}$. The matrix $\bm{B}=(B_{ij})_{1\leq i,j\leq n}\in\mathbb{N}_{0}^{n\times n}$ represents the bonds and their orders between atoms, i.e., we have that $B_{ij}\in\mathbb{N}_{0}$ is the bond order between the $i$-th and the $j$-th atom. In particular, $\bm{B}$ is symmetric with zero diagonal and nonzero columns and rows (i.e., each atom is connected to at least one other atom and not to itself). For example, given the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$ the molecule with SMILES representation C=O can be given as $(\bm{a},\bm{B})$ with $\bm{a}=(1,3)$ and $\bm{B}=\begin{pmatrix}0&2\\ 2&0\end{pmatrix}$.
We denote by $\mathcal{M^{\prime}}$ the space of all molecular graphs $m=(\bm{a},\bm{B})$ as described above. Accordingly, let $\mathcal{M}\subset\mathcal{M^{\prime}}$ be the subspace of molecular graphs that are chemically valid, i.e., all non-hydrogen atoms follow the octet rule. We refer to following the octet rule as satisfying the ’valence constraints’ throughout the paper.
Sequential molecular graph design
Similar to Zhou et al. (2019), we pose molecular design as a sequential Markov decision process (MDP), where an agent assembles a molecular graph by iteratively adding atoms or increasing the bond order between atoms. We note that in the graph, hydrogens are always only considered implicitly, and an addition of an atom or a bond leads to a replacement of an implicit hydrogen.
On a high level, a single molecule is constructed by the agent as follows: The agent observes an initial molecule $m^{(0)}$, and then performs some action $x^{(0)}$ on it, which results in a new molecule $m^{(1)}$. Then, the agent observes $m^{(1)}$, decides on an action $x^{(1)}$ leading to molecule $m^{(2)}$, and so on. This iterative design process terminates at some molecule $m^{(T)}$ once the agent decides to perform a special DontChange action, which does not alter the molecular graph but rather marks the design as finished.
We formalize this as follows: Let $m^{(t)}=(\bm{a}^{(t)},\bm{B}^{(t)})$ be some molecule with atoms $\bm{a}^{(t)}\in\{1,\dots,k\}^{n^{(t)}}$ and bond matrix $\bm{B}^{(t)}\in\mathbb{N}_{0}^{n^{(t)}\times n^{(t)}}$ as above. We transition to a new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$ by performing some action $x^{(t)}\in\mathcal{A}$ on $m^{(t)}$. An action $x^{(t)}$ in the action space $\mathcal{A}$ is of one of the following three types:
  1. 1.
    $x^{(t)}=\textsc{DontChange}$, which terminates the design of the molecule. In particular, $m^{(t+1)}=m^{(t)}$ and $n^{(t+1)}=n^{(t)}$.
  2. 2.
    $x^{(t)}=\textsc{AddAtom}(j,l)$, with $j\in\{1,\dots,k\}$ and $l\in\{1,\dots,n^{(t)}\}$. This action adds an atom of type $\Sigma_{j}$ to the graph and bonds it to the $l$-th atom. In particular, we set $n^{(t+1)}=n^{(t)}+1$. For the new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$, the atom vector $\bm{a}^{(t+1)}\in\mathbb{N}^{n^{(t+1)}}$ is obtained by appending $j$ to $\bm{a}^{(t)}$. The second entry $l\in\{1,\dots,n^{(t)}\}$ indicates that we bond this new $(n+1)$-th atom to the $l$-th atom, i.e., $\bm{B}^{(t+1)}\in\mathbb{N}_{0}^{n^{(t+1)}\times n^{(t+1)}}$ is obtained by appending a zero row and column to $\bm{B}^{(t)}$ and setting $B^{(t+1)}_{n+1,l}=B^{(t+1)}_{l,n+1}=1$.
  3. 3.
    $x^{(t)}=\textsc{IncBond}(j,l)$, meaning that we are increasing the bond order between atoms $j$ and $l$. In particular, we have $n^{(t+1)}=n^{(t)}$ and $\bm{a}^{(t+1)}=\bm{a}^{(t)}$. The bond matrix $\bm{B}^{(t+1)}$ is obtained from $\bm{B}^{(t)}$ by setting $B^{(t+1)}_{j,l}=B^{(t+1)}_{l,j}=B^{(t)}_{j,l}+1$. Note that this allows not only higher order bonds, but also bonds between previously unbonded atoms.
Given a molecule $m^{(0)}$ and a sequence of actions $x^{(0)},\dots,x^{(t-1)}$, we will also write $(m^{(0)},x^{(0)},\dots,x^{(t-1)})$ for the molecule $m^{(t)}\in\mathcal{M}$ that results from $m^{(0)}$ by sequentially applying the actions $x^{(0)},\dots,x^{(t-1)}$ to $m^{(0)}$. We note that, in general, the action sequence $x^{(0)},\dots,x^{(t-1)}$ is not unique to from $m^{(0)}$ to $m^{(t)}$.
Starting from a molecule in $\mathcal{M}$, by removing chemically invalid actions from the action space (which would lead to violation of valence constraints) at each step, we can guarantee to stay in the space of chemically valid molecules. Once the agent chooses the DontChange action, the design process is considered finished. We note that, starting from an appropriate initial atom, it is in fact possible to reach every target molecule in the chemically valid space $\mathcal{M}$ (i.e., all molecular graphs that can be constructed from the alphabet $\Sigma$ and that satisfy the valence constraints) when starting from any atom that exists in the target molecule. For an illustrative molecule construction, see Fig. 1.
Each addition of a bond or atom removes an implicit hydrogen from the respective atom, allowing for a straightforward representation of ionization states: Ions can be incorporated simply by adding charged atom types to the predefined alphabet, without the need for additional modifications to the action space. In contrast, the graph-based approach can not handle stereochemistry in a straightforward way.
Figure 1: Example for the sequential application of actions $x^{(0)},\dots,x^{(3)}$ to a molecule $m^{(0)}$, using the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$. We show the index of each atom, which can be arbitrarily chosen at the beginning. Light blue indicates where in the graph an action is applied. The last action is DontChange, which does not change the molecular graph, but marks it as a finished design.
Molecular optimization
Formally, we aim to design molecules
$m^{*}\in\operatorname*{arg\,max}_{m\in\mathcal{M}}f(m)$ (1)
that maximize a predefined objective function $f\colon\mathcal{M}^{\prime}\to\mathbb{R}\cup\{-\infty\}$, where chemically invalid molecules are mapped to $\infty$- ∞. As in previous work (Mazuz et al., 2023; Zhou et al., 2019; Xu et al., 2024), we pose the corresponding learning problem to (1) in the terms of deep RL: the agent’s decision at each step is guided by a policy that maps a chemically valid molecule to a probability distribution over possible actions. The policy is modeled by a neural network: We write $\pi_{\theta}\colon\mathcal{M}\to\Delta\mathcal{A}$ for a policy depending on network parameters $\theta$, that maps a valid molecule $m\in\mathcal{M}$ to a probability distribution $\pi_{\theta}(m)$ over $\mathcal{A}$. The goal is to find $\pi_{\theta}$ that, given any initial molecule $m^{(0)}$, maximizes the expectation
$\mathbb{E}_{\begin{subarray}{c}x^{(0)},\dots,x^{(T)}\\ x^{(T)}=\textsc{DontChange}\end{subarray}\sim\pi_{\theta}}\left[f\left((m^{(0)},x^{(0)},\dots,x^{(T)})\right)\right],$ (2)
where the expectation is taken over finished molecules sampled from $\pi_{\theta}$.
Satisfying constraints
To ensure chemical validity at every step of the molecule design process, we mask any action in the policy that would lead to a violation of valence constraints. That means, that we set the corresponding probability to zero in the distribution predicted by the policy (before re-normalizing the distribution).
Our graph-based approach allows us to extend this concept of action masking to enforce additional constraints, such as particular bonding patterns, minimum/maximum number of atoms and their types, or the presence of structural motifs like rings. We explore these constraints in detail in Section 3. We note that it is possible to simply assign an objective value of $\infty$- ∞ to molecules that violate these constraints after the design process. However, the ability to preemptively avoid constraint-violating regions by masking actions during the design process can significantly reduce the search space. In contrast, methods that generate molecules by sequentially synthesizing SMILES (Gupta et al., 2018; Xu et al., 2024; Mazuz et al., 2023) strings struggle to enforce such constraints during the construction process while also adhering to the grammatical rules of SMILES.
2.2 Learning algorithm
We now introduce our method for training the policy neural network $\pi_{\theta}$. Our proposed learning algorithm is a hybrid of the deep cross-entropy method (CEM) (Wagner, 2021) and the self-improvement learning (SIL) (Huang et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024a, b) strategy. Both approaches train the network over multiple epochs in a self-improving loop, where, in each epoch, the current policy is used to generate a set of action sequences. The best sequences are selected as ’pseudo-labels’, i.e., they serve as the dataset for supervised training. The network is then trained for one epoch to assign higher probabilities to these sequences before the process is repeated with the updated network to generate new action sequences. Unlike other deep RL algorithms, this approach does not require reward shaping or value approximation. Furthermore, training in a supervised way provides stability and allows for the use of larger, decoder-only transformer architectures (Vaswani, 2017), which would be prohibitively slow under standard RL. However, SIL and the deep CEM differ in key aspects: The deep CEM is formulated for problems with a single instance (as in our molecular design task) and retains a certain percentage of the best action sequences in each epoch, which are obtained through simple sampling from the policy’s predicted distributions at each step. SIL, on the other hand, is typically applied to problems with infinitely many instances and employs more advanced sequence decoding techniques, such as sequence sampling without replacement (Kool et al., 2019), to improve solution quality and diversity while maintaining efficient decoding speed.
In our proposed approach, we aim to develop a method that works for single-instance problems, as in the deep CEM, while retaining the diverse sampling mechanism used in SIL methods (Pirnay & Grimm, 2024a, b). The pseudocode for our approach is presented in Algorithm 1, and we describe the key steps below:
  1. 1.
    We begin with a policy network $\pi_{\theta}$. The parameters $\theta$ can be initialized randomly or, e.g., pretrained in an unsupervised manner on existing molecules (see Section 3.3). Additionally, we assume an initial molecule $m_{0}\in\mathcal{M}$ in the space of chemically valid molecules $\mathcal{M}$, from which the construction starts. In practice, we typically choose $m_{0}$ to consist of a single carbon atom.
  2. 2.
    Throughout the training process, we maintain a set BestFound of the best molecules discovered, which will be returned at the end. Initially, this set contains only the starting molecule $m_{0}$.
  3. 3.
    In each epoch, we incrementally sample action sequences without replacement from the policy until either an improved molecule is found or a maximum number of samples is reached (lines 5-10). For sampling sequences without replacement, we use the efficient batch-wise incremental method introduced by Shi et al. (2020), which relates to Stochastic Beam Search (Kool et al., 2019) and is employed in SIL methods (Pirnay & Grimm, 2024a, b). Sampling without replacement ensures that unique action sequences are generated, exploring the search space effectively (especially for shorter action sequences (Shi et al., 2020)). We emphasize that, while different action sequences may lead to the same molecule, this does not pose an issue because the policy is only concerned with generating the action sequences rather than the molecular structure. Since the goal of sampling sequences without replacement is to produce distinct sequences, the policy still generates diverse outputs, even if the resulting molecules can be identical.
  4. 4.
    After sampling, the set BestFound serves as a training dataset (lines 12-14). Similar to how decoder-only models in language modeling are trained to predict the next token from partial text, we sample batches of intermediate molecules and train the network with a cross-entropy loss to predict the next action for the corresponding molecule.
  5. 5.
    In the next epoch, the process is repeated with the updated network weights.
Input: $\pi_{\theta}$: Policy network with trainable parameters $\theta$
Input: $f$: objective function to maximize
Input: $m_{0}$: initial molecule
Input: $s\in\mathbb{N}$: number of best molecules to keep
Input: $q_{\min},q_{\max}\in\mathbb{N}$: minimum and maximum number of molecules to sample without replacement
1 $\textsc{BestFound}\leftarrow\{m_{0}\}$
2 $\textsc{BestObj}\leftarrow f(m_{0})$
3 foreach epoch do
4       $\textsc{Sampled}\leftarrow\emptyset$
5       while $|\textsc{Sampled}|<q_{\max}$ do
6             $m\leftarrow$ sample from $\mathcal{M}$ with probability $\pi_{\theta}\left(m\ |\ m_{0},m\notin\textsc{Sampled}\right)$
7             $\textsc{Sampled}\leftarrow\textsc{Sampled}\cup\{m\}$
8             if $f(m)>\textsc{BestObj}$ then
9                   $\textsc{BestObj}\leftarrow f(m)$
10                   if $|\textsc{Sampled}|>q_{\min}$ then
11                         break
12
13
[…]
15       foreach batch do
16             uniformly sample from BestFound batch of $B$ (intermediate) molecules $m_{t_{j+1}}^{(j)}=(m_{t_{j}}^{(j)},x_{t_{j}}^{(j)})$ for $j\in\{1,\dots,B\}$
17             minimize batch-wise cross-entropy loss $L_{\theta}=-\frac{1}{B}\sum_{j=1}^{B}\log\pi_{\theta}\left(x_{t_{j}}^{(j)}\ |\ m_{t_{j}}^{(j)}\right)$
18
19
return BestFound
Algorithm 1 Learning algorithm for molecular design
2.3 Policy network architecture
The policy network takes a molecule as input and predicts a probability distribution over next actions. To achieve this, we use a simplified version of the Graphormer (Ying et al., 2021) architecture, which treats the molecule’s atoms as an unordered sequence of nodes. This sequence is passed through a stack of transformer layers, with the attention mechanism augmented by bonding information. The resulting latent representations of the nodes are then used to predict action distributions. Figure 2 provides an illustration, and we elaborate on the architecture below:
Figure 2: Flow of a molecule through the policy network of our method GraphXForm. a. We consider the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$. The molecule’s underlying graph is augmented with a virtual node (indexed with 0) and embedded into the latent space $\mathbb{R}^{d}$. b. The latent sequence of atoms is passed through a stack of ReZero transformer layers, omitting positional encoding. In the multi-head attention, individual attention scores between atoms are biased with learnable scalars that depend on their bond order. These bias terms are learnable for each transformer layer and attention head individually. c. The sequence output by the transformer is projected through linear layers to generate logits for the distributions $P^{(0)},P^{(1)}$ and $P^{(2)}$.
Molecular graph transformer
Let $m=(\bm{a},\bm{B})$ represent a molecule, where $\bm{a}=(a_{1},\dots,a_{n})\in\mathbb{N}^{n}$ are the atoms, and $\bm{B}\in\mathbb{N}_{0}^{n\times n}$ is the bond matrix. Each atom $a_{i}\in\{1,\dots,k\}$ corresponds to an atom type selected from an alphabet $\Sigma=(\Sigma_{1},\dots,\Sigma_{k})$.
Following the Graphormer (Ying et al., 2021) model, we introduce a ‘virtual atom’ $a_{0}=0$ into the molecular graph, which is connected to every atom via a special ‘virtual bond’ (see Figure 2). This virtual bond is represented in the bond matrix by an integer outside the range of standard bond orders. The virtual atom functions similarly to the special [CLS]-token in BERT (Devlin, 2018) and accumulates sequence-level information for downstream tasks. Conceptually, the virtual atom acts as an additional message proxy between nodes in the molecular graph.
The sequence $(a_{0},a_{1},\dots,a_{n})$ is embedded into a corresponding sequence of latent node representations $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1},\dots,\bm{\hat{a}}_{n})\in\mathbb{R}^{d}$, where each $\bm{\hat{a}}_{i}$ is a learnable vector associated with the atom type.
Additionally, let $\bm{z}_{i}\in\mathbb{R}^{d}$ be a learnable embedding that depends on the number of bonds formed by the $i$-th atom. The augmented sequence $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1}+\bm{z}_{1},\dots,\bm{\hat{a}}_{n}+\bm{z}_{n})$ is then passed through a stack of transformer layers using ReZero normalization (Bachlechner et al., 2021). Importantly, to preserve the permutation equivariance of the network (i.e., the order of the atoms does not matter), no positional encoding is applied.
To incorporate bonding information within the transformer layers, let $(\bm{A}_{ij})_{0\leq i,j\leq n}\in\mathbb{R}^{(n+1)\times(n+1)}$ be the computed self-attention matrix in a layer (for any attention head) corresponding to the input sequence $(a_{0},a_{1},\dots,a_{n})$. As in Graphormer (Ying et al., 2021), and similar to the Molecule Attention Transformer (Maziarka et al., 2019), we augment the attention matrix $\bm{A}$ by introducing bond-specific information. This is done by adding a learnable bias to the attention scores before applying the softmax operation. Specifically, for $0\leq i,j\leq n$, the attention score $\bm{A}_{ij}$ between the $i$-th and $j$-th sequence elements is modified as follows:
$\bm{A}_{ij}\leftarrow\bm{A}_{ij}+b_{ij},$ (3)
where $b_{ij}\in\mathbb{R}$ is a learnable scalar that only depends on the bond order $B_{ij}$ for $i,j>0$ (and not on the indices $i,j$). In particular, for the special bonds involving the virtual atom (i.e., if $i=0$ or $j=0$), $b_{ij}=b$ is a learnable scalar shared across all atoms. We note that $b_{ij}$ for each bond order is learned independently across layers and attention heads.
Action distribution
The stack of transformer layers produces a transformed sequence of node embeddings $(\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n})$, where each $\bm{e}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th atom $a_{i}$ in the original sequence. This sequence is used to predict three distributions, $P^{(0)},P^{(1)},P^{(2)}$, which together factorize the full distribution over possible next actions.
Informally, $P^{(0)}$ represents a distribution over whether to terminate the synthesis, add a new atom, or increase the bond order between two existing atoms. The distribution $P^{(1)}$ specifies which atom to select, either as the atom to bond a newly added atom to, or as the first atom of a pair for increasing bond order. Finally, $P^{(2)}$ is used in the case of bond increase to specify the second atom in the pair. Formally, the unnormalized log-probabilities (logits) for these distributions are obtained as follows:
  • $P^{(0)}$: We compute the logits for $P^{(0)}=(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k},p^{(0)}_{k+1})$ by projecting the vector $\bm{e}_{0}$ using a linear layer $g_{0}\colon\mathbb{R}^{d}\to\mathbb{R}^{k+2}$. Here, $p^{(0)}_{0}$ is the probability of choosing the action DontChange. For $1\leq j\leq k$, $p_{j}$ represents the probability of adding an atom of type $\Sigma_{k}$ (i.e., $\textsc{AddAtom}(j,\cdot)$ without specifying the bonding atom yet). Lastly, $p^{(0)}_{k+1}$ is the probability of increasing the bond between two atoms ($\textsc{IncBond}(\cdot,\cdot)$) without specifying the atom pair.
  • $P^{(1)}$: To obtain $P^{(1)}=(p^{(1)}_{1},\dots,p^{(1)}_{n})$, we apply a linear layer $g_{1}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each $\bm{e}_{1},\dots,\bm{e}_{n}$. Then, $p^{(1)}_{j}$ represents the probability of selecting the $j$-th atom to bond to in $\textsc{AddAtom}(\cdot,\cdot)$, or as the first atom in $\textsc{IncBond}(\cdot,\cdot)$.
  • $P^{(2)}$: Similarly, for $P^{(2)}=(p^{(2)}_{1},\dots,p^{(2)}_{n})$, we apply another linear layer $g_{2}\colon\mathbb{R}^{d}\to\mathbb{R}$ to compute the logits for the probability $p^{(2)}_{l}$, representing the likelihood of selecting the $l$-th atom as the second atom in IncBond.
In practice, during inference, we use $P^{(0)}$ and $P^{(1)}$ directly to choose the first atom for both AddAtom and IncBond, and discard $P^{(2)}$. If IncBond has been selected from $P^{(0)}$, we reprocess the molecule through the network, marking the atom chosen from $P^{(1)}$ by adding a distinct learnable embedding to its corresponding sequence element. In this second pass, we use only $P^{(2)}$ to select the second atom for IncBond.
This multi-step action prediction allows us to easily mask actions (i.e., setting their probability to zero in the policy) that would lead to invalid molecules. While checking for actions that would violate constraints adds a small amount of computational overhead, it has a significant benefit: invalid molecules can be immediately disregarded, preventing the network from wasting resources on infeasible designs. Masking invalid actions not only reduces the search space but also speeds up training by avoiding the need for the model to learn through trial and error how to avoid invalid molecules.
3 Results and Discussion
First, we outline our design tasks, including the objective functions and property prediction methods. Next, we describe the experimental setup common to all methods and outline the specific setup and hyperparameters for our method, followed by a brief overview of the methods used for comparison. We then present preliminary baseline objectives obtained from screening a list of available molecules and discuss our results in relation to those of the other methods.
3.1 Example solvent design tasks
Objectives
We aim to design solvents for two-phase aqueous-organic systems used in liquid-liquid extraction. Our focus is on two examples motivated by biotechnology, where products are typically produced in an aqueous solution using microorganisms or enzymes. In such processes, products are to be extracted using the organic solvents we aim to design. We assume a spatially uniform temperature of 298 K in both examples.
The first task focuses on the separation of isobutanol (IBA) from water, a common liquid-liquid extraction process. The chosen solvent should be largely immiscible with water (i.e., low solubility exhibited for both the solvent in water and water in the solvent) and possess high affinity for IBA. As is common practice in chemical engineering, we use the partition coefficient $P_{\text{IBA}}^{\infty}$ at small mole fractions of IBA in both phases $x_{\text{IBA}}$:
$P_{\text{IBA}}^{\infty}=\lim_{x_{\text{IBA}}^{\text{W}}\to 0}\frac{x_{\text{IBA}}^{\text{S}}}{x_{\text{IBA}}^{\text{W}}}$ (4)
where $x_{\text{IBA}}^{\text{W}}$ and $x_{\text{IBA}}^{\text{S}}$ are the mole fractions of IBA in water (W) and the solvent (S), respectively. This coefficient serves as a simple yet effective measure of the relative affinity of the solvent compared to water. Assuming low mutual solubility between the solvent and water, $P_{\text{IBA}}^{\infty}$ can be well approximated by the ratio of IBA’s activity coefficients at infinite dilution, $\gamma_{\text{IBA,W}}^{\infty}$ and $\gamma_{\text{IBA,S}}^{\infty}$, in water and solvent, respectively:
$P_{\text{IBA}}^{\infty}=\frac{\gamma_{\text{IBA,W}}^{\infty}}{\gamma_{\text{IBA,S}}^{\infty}}.$ (5)
To ensure the formation of two phases, i.e., a miscibility gap between the solvent and water, we use the following constraint:
$\gamma_{\text{S,W}}^{\infty}\cdot\gamma_{\text{W,S}}^{\infty}>\exp(4).$ (6)
This constraint guarantees a phase split between the water and solvent, assuming that the activity coefficient profiles follow the two-parameter Margules $g^{E}$ model (Wisniak, 1983). Although the activity coefficients of all conceivable solvent/water mixtures will not necessarily follow this model, the constraint still serves as a useful indicator for miscibility gaps.
The partition coefficient and the miscibility gap constraint are then combined to form the following scalar objective function:
$\max\frac{1}{\gamma_{\text{IBA,S}}^{\infty}}+\left(\tanh\left(\gamma_{\text{S,W}}^{\infty}\cdot\gamma_{\text{W,S}}^{\infty}-\exp(4)\right)-1\right)\cdot 10.$ (7)
Hereby, the solvent-independent constant $\gamma_{\text{IBA,W}}^{\infty}$ is omitted.
Our second task centers on an extraction process presented by Peters et al. (2008), who carried out a solvent screening using COSMO-RS as a property predictor. Here, an enzymatic reaction in aqueous medium converts 3,5-dimethoxy-benzaldehyde (DMBA) molecules to (R)-3,3’,5,5’-tetra-methoxy-benzoin (TMB). The task is to find an organic solvent that forms a two-phase system with water. Similar to the IBA task, an optimal solvent should have a high affinity for the product TMB, enabling it to pull TMB out of the aqueous phase. At the same time, however, the solvent should have a low affinity for the educt DMBA. Designing a suitable solvent for this task is extremely challenging because DMBA and TMB possess similar chemical structures and polarities.
Again assuming small concentrations of DMBA and TMB as well as low mutual solubility between the solvent and water, we define the following partition coefficients similarly to our IBA task:
$P_{\text{DMBA}}^{\infty}=\frac{\gamma_{\text{DMBA,W}}^{\infty}}{\gamma_{\text{DMBA,S}}^{\infty}}$ (8)
$P_{\text{TMB}}^{\infty}=\frac{\gamma_{\text{TMB,W}}^{\infty}}{\gamma_{\text{TMB,S}}^{\infty}}.$
Following Peters et al. (2008), we maximize the ratio $P_{\text{TMB}}^{\infty}/P_{\text{DMBA}}^{\infty}$, while additionally enforcing the miscibility gap constraint from Equation 6 leading to the following scalar objective:
$\max\frac{\gamma_{\text{TMB,S}}^{\infty}}{\gamma_{\text{DBMA,S}}^{\infty}}+\left(\tanh\left(\gamma_{\text{S,W}}^{\infty}\cdot\gamma_{\text{W,S}}^{\infty}-\exp(4)\right)-1\right)\cdot 10.$ (9)
Hereby, we again omitted the constants $\gamma_{\text{TMB,W}}^{\infty}$ and $\gamma_{\text{DBMA,W}}^{\infty}$.
Property prediction
To obtain activity coefficients at infinite dilution, we use a state-of-the-art graph neural network (GNN)  (Gilmer et al., 2017; Reiser et al., 2022; Rittig et al., 2023c). Specifically, we employ the Gibbs-Helmholtz (GH-) GNN that was developed by Medina et al. (2023) for predicting activity coefficients at infinite dilution of binary mixtures at varying temperature. The GH-GNN takes the molecular graphs of the two molecules within a binary mixture and the temperature as inputs. First, structural information from the individual molecular graphs and molecular (self-)interactions based on a mixture graph are encoded into a vector representation, the mixture fingerprint. Based on the mixture fingerprint, a multilayer perceptron (MLP) then predicts the parameters of the Gibbs-Helmholtz relationship so that infinite dilution activity coefficients can be predicted with temperature. The GH-GNN is trained in a structure-to-property manner, i.e., it directly learns activity coefficients at infinite dilution as a function of the molecular graphs. A large data set of experimental activity coefficients at infinite dilution from the DECHEMA Chemistry Data Series  (Gmehling et al., 2008) was used for training. For further details on the GH-GNN architecture, we refer to Medina et al. (2023).
We note that many GNN models and other ML models such as transformers have been developed for activity coefficient prediction  (Rittig et al., 2023a; Winter et al., 2022; Damay et al., 2021), also considering the composition-dependency and thermodynamic consistency  (Rittig et al., 2023b; Winter et al., 2023; Rittig & Mitsos, 2024; Specht et al., 2024). We here chose the GH-GNN as it is specialized for activity coefficients at infinite dilution and was trained on a much larger experimental database than the other models, thus covering a larger chemical space  (Medina et al., 2023), which is desirable for molecular design. This model has shown high prediction accuracy, outperforming well-established methods for predicting activity coefficients at infinite dilution such as UNIFAC  (Fredenslund et al., 1975) or COSMO-RS  (Klamt et al., 2010), cf.  (Medina et al., 2023). Future work could investigate further additional activity coefficient models or directly predicting partition coefficients with ML  (Zamora et al., 2023; Nevolianis et al., 2024).
3.2 General computational setup
To ensure comparability across methods in the design task, we restrict the set of available atoms to carbon, nitrogen, and oxygen (alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$). In the solvent design tasks we consider, molecules are evaluated using objective functions based on a surrogate model for predicting activity coefficients (see Section 3.1). To avoid potential exploitation of the surrogate model by constructing large molecules that are beyond the size of typically used solvents and thus likely to be different from the molecules used to train the surrogate model, we constrain the design to molecules of a moderate size, with a maximum of 25 atoms. The runtime of a method can vary significantly due to differences in implementation or available hardware. One approach to align computational resources is to limit the number of calls to the objective function (Gao et al., 2022). While this method is useful for comparing sample efficiency, it provides limited insight into the overall efficiency of a method when objective function evaluations are relatively inexpensive. Given the efficiency of the surrogate model, objective function evaluations can be parallelized and run in batches very quickly. Therefore, we take a practical approach to align computational resources: each design run of a method is allocated a time budget of eight wall-clock hours on identical hardware, utilizing a single NVIDIA H100 GPU with 80 GB of memory.
3.3 Our method: GraphXForm
Network hyperparameters
For all our experiments, we set the dimension of the latent space $\mathbb{R}^{d}$ to $d=128$ (see Section 2.3). The network consists of eight transformer layers with ReZero normalization, each with eight heads in the multi-head attention, and a feed-forward dimension of 512.
Pretraining
While it is possible for the agent to learn without any prior knowledge of existing compounds, we pretrain the network on known molecules using SMILES strings from the ChEMBL database (Davies et al., 2015). We filtered the database to ensure the molecules contain only atoms within the alphabet $\Sigma$, yielding approximately 370,000 molecules. Each SMILES string is then converted into an action sequence in our graph formulation. Since this conversion is not unique, we arbitrarily select one possible action sequence. The resulting sequences are used to train the model in a self-supervised way, predicting the next action given the previous actions, as outlined in lines 12-14 in Algorithm 1. Training is conducted with a batch size of 128, over a total of 2.5 million batches.
Petraining the network on existing molecules establishes a prior that captures the characteristics of real molecules, which is crucial for most generative methods such as REINVENT (Olivecrona et al., 2017; Xu et al., 2024) or JT-VAE (Jin et al., 2018).
Fine-tuning
Given an objective function $f\colon\mathcal{M}^{\prime}\to\mathbb{R}\cup\{-\infty\}$, we fine-tune the pretrained policy network using the learning algorithm described in Section 2.2. Unless stated otherwise, we initialize the molecule $m_{0}$ as a single carbon atom. We set $s=500$ as the number of top molecules to retain throughout training. In each epoch, we sample between $q_{\text{min}}=10\cdot 1024$ and maximum of $q_{\text{max}}=50\cdot 1024$ unique action sequences. At the end of each epoch, we train the network on 20 batches of size 64, sampled uniformly from the top 500 molecules. We intentionally keep the number of training batches per epoch relatively low for the network’s size to prevent premature overfitting in all considered cases. However, we observe that in most runs, training on more batches does not adversely affect performance and can actually increase the speed of convergence.
3.4 Benchmark methods
We compare our proposed approach with four diverse molecular design methods: Graph GA (Jensen, 2019) (a genetic algorithm (GA) applied to molecular graphs), REINVENT-Transformer (Xu et al., 2024) (a transformer model that generates SMILES strings), Junction Tree VAE (Jin et al., 2018) combined with a GA for optimization in latent molecular space, and STONED (Nigam et al., 2021) (which performs string mutations on the SELFIES (Krenn et al., 2020) representation). For benchmarking, we use the source code provided by Gao et al. (2022). Although these methods employ different paradigms, we apply the same objective function across all methods to guide the design process.
Graph GA and STONED
Graph GA (Jensen, 2019) uses a GA that directly operates on the molecular graph by mutating atoms and fragments, using crossover rules derived from graph matching. It is a strong non-learning method that has been shown to outperform several SMILES-based learning approaches (Jensen, 2019; Gao et al., 2022). In contrast, STONED (Nigam et al., 2021) employs a GA that operates on the string level and manipulates tokens within the SELFIES molecular representation (Krenn et al., 2020).
REINVENT-Transformer
REINVENT-Transformer (Xu et al., 2024) is a recent state-of-the-art method for designing molecules by synthesizing SMILES strings in an autoregressive way. The approach involves a transformer network that is pretrained in a self-supervised manner on known molecules and then fine-tuned with reinforcement learning using a variant of the REINFORCE algorithm (Williams, 1992). REINVENT-Transformer is similar to GraphXForm in that both rely on pretrained transformers and generate molecules autoregressively. However, REINVENT-Transformer constructs molecules in a language-like manner by building a string token by token from an allowed vocabulary, whereas GraphXForm operates directly on the molecular graph by adding atoms and bonds.
This direct manipulation of the molecular graph allows GraphXForm to guarantee the design of feasible molecules by construction, whereas REINVENT-Transformer must learn the SMILES grammar to avoid invalid structures. Additionally, our approach can start from arbitrary structures, unlike the left-to-right string construction in REINVENT-Transformer, making it straightforward to encode specific constraints to limit the search space. For a fair comparison, we pretrain the policy for REINVENT-Transformer using the same dataset as GraphXForm.
Junction Tree VAE
We further use variational autoencoders (VAEs) that have been widely used as a generative model in ML-guided molecular design, cf. overviews in Refs.  (Sanchez-Lengeling & Aspuru-Guzik, 2018; Bilodeau et al., 2022; Anstine & Isayev, 2023). VAEs use an encoder-decoder structure: The encoder maps molecules to a continuous vector representation, referred to as latent space, from which the decoder reconstructs these molecules. The molecular latent space can facilitate exploration of the molecular space. That is, different optimization strategies can be employed to discover points in the latent space that correspond to promising novel molecules. These strategies include random sampling, Bayesian optimization, and GAs (Sanchez-Lengeling & Aspuru-Guzik, 2018; Rittig et al., 2023d; Anstine & Isayev, 2023).
We herein use the Junction Tree VAE (JT-VAE)  (Jin et al., 2018) in combination with GAs. The JT-VAE operates on molecular graphs and their non-cyclic abstractions (junction trees), and has shown a high rate of decoding vectors from its latent space to chemically valid molecules. For the solvent design task, we only consider molecules that conform to the alphabet $\Sigma$. We therefore train JT-VAE on a subset of the QM9 data set, which consists of about 128,000 molecules containing at most nine heavy atoms. We use GAs to optimize within the JT-VAE latent space because they enable exploring a large number of latent vectors, hence molecules, due to the linear scaling with the number of sampling points.
3.5 Design results
The goal for all methods is to find suitable solvents with respect to the two objectives from Equations 7 and 9 for the two example problems, using the general setup outlined in Section 3.2.
Table 1: Performance of different molecular design methods for the two example tasks. Each method is run across three different random seeds, with a time budget of 8 hours. We report the objective function evaluation of the best molecule found over all runs (‘max best’), best averaged over all the three runs $\pm$ standard deviation (‘mean best’), the average of the top 20 molecules of the best run (‘max top 20’), and the mean of the top 20 over all three runs $\pm$ standard deviation (‘mean top 20’).
Method IBA (cf. (7)) TMB/DMBA (cf. (9))
max best mean best max top 20 mean top 20 max best mean best max top 20 mean top 20
JT-VAE (Jin et al., 2018) 6.85 6.41 $\pm$ 0.66 6.04 5.68 $\pm$ 0.57 2.16 1.56 $\pm$ 0.54 1.44 1.20 $\pm$ 0.23
STONED (Nigam et al., 2021) 8.31 7.42 $\pm$ 0.94 6.72 6.28 $\pm$ 0.41 2.39 1.68 $\pm$ 0.65 1.91 1.45 $\pm$ 0.49
Graph GA (Jensen, 2019) 8.87 7.13 $\pm$ 3.01 8.67 6.80 $\pm$ 3.22 8.40 8.14 $\pm$ 0.27 8.07 7.95 $\pm$ 0.32
REINVENT-Transformer (Xu et al., 2024) 8.87 8.32 $\pm$ 0.52 8.66 8.09 $\pm$ 0.45 7.41 6.52 $\pm$ 1.17 7.22 6.42 $\pm$ 0.96
GraphXForm (ours) 8.87 8.87 $\pm$ 0.00 8.67 8.67 $\pm$ 0.00 9.41 9.33 $\pm$ 0.12 9.23 9.14 $\pm$ 0.11
Screening list
To contextualize the results of all methods, we selected all 6,098 compounds from the COSMObase 2020 database that conform to the alphabet $\Sigma$. We calculated the objective functions for those compounds that also meet the miscibility gap requirements, cf. (6). The top three compounds based on their objective values are shown in the first column in Figures 3 and 4. For the IBA task, the best molecule in this list achieves an objective value of 5.57, while for the TMB/DMBA task, the highest objective value is 3.03.
Model results
Table 1 compares the performance of the different molecular design methods for the two solvent design tasks. We report the objective value of the best molecule found by each method, as well as the average objective values of the top 20 molecules. We also present results averaged over multiple runs with different random seeds, providing insight into the robustness of each method. Additionally, Figures 3 and 4 display the structural formulas of the top three molecules identified by each method.
Figure 3: IBA task: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
Figure 4: TMB/DMBA task: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
In the IBA task, GraphXForm, REINVENT-Transformer, and Graph GA all identified the same best molecule, which has an objective value of 8.87. However, GraphXForm consistently found this molecule in every run, highlighting its stability. For example, in contrast, while Graph GA achieved the same average value for the top 20 molecules, its mean best objective was only 7.13.
For the more challenging TMB/DMBA task, GraphXForm significantly outperforms all others across every metric. Notably, Graph GA performs better overall than the recent REINVENT-Transformer, likely due to REINVENT-Transformer’s sensitivity to initialization during RL fine-tuning. This is also reflected in its relatively high standard deviation compared to the other methods.
Interestingly, the JT-VAE and STONED methods identified molecules with significantly lower objective values for the TMB/DMBA task when compared to other methods, with maximum values of approximately 2.16 and 2.39, respectively. However, their results for the IBA task (6.85 and 7.53) were closer to those of the other methods. We attribute this discrepancy to the nature of the tasks: The TMB/DMBA task is inherently more challenging and requires larger, more complex molecules, while the best-performing molecules in the IBA task were smaller. As JT-VAE was trained only on molecules with up to nine heavy atoms, its ability to explore larger molecules seems limited. Similarly, STONED struggled to effectively explore larger molecular structures.
In summary, GraphXForm demonstrates highly promising results in both solvent design tasks, outperforming state-of-the-art molecular design methods in terms of identifying the best candidates and ensuring robustness in the design process. In the following sections, we further explore the flexibility of GraphXForm by imposing structural constraints on the designed molecules and starting the design process from initial molecular structures.
3.6 Including structural constraints in molecular design
In addition to optimizing an objective, such as a physical property, it is often desirable to enforce specific structural characteristics in the generated molecular candidates. These structural constraints can include limitations on ring size or specific bonding patterns between atom types, which can improve chemical feasibility, such as enhancing synthesizability and stability or reducing toxic moieties. For instance, in the TMB/DMBA task, although the top molecules generated by GraphXForm exhibit high objective values, they possess several characteristics that pose challenges for solvent design. Notably, the presence of single nitrogen-nitrogen bonds often leads to instability (hyd, 2024), resulting in reactive compounds that are unsuitable for liquid-liquid extraction. Additionally, the existence of a nine-member ring could pose significant challenges to the solvent’s synthesizability (Steinborn et al., 2023).
As noted in Section 2.1, string-based molecular design methods cannot directly enforce such structural constraints. Typically, this is managed by assigning a very low objective value to molecules that violate these constraints. In contrast, by operating directly on the molecular graph, GraphXForm allows us to restrict the search space for many constraints by simply masking actions that would lead to constraint-violating molecules. To examine how structural constraints affect the performance of GraphXForm in finding molecules with high objective values, we impose the following constraints: We limit the ring size to a maximum of six atoms and disallow any bonds between two oxygen atoms as well as single bonds between two nitrogen atoms. We note that these constraints of course do not cover all possibilities for maintaining synthetic accessibility and chemical stability; instead, they serve as examples to illustrate the capabilities of GraphXForm.
For the IBA task, the best molecule identified in Table 1 with objective value of 8.87 contains multiple nitrogen-nitrogen single bonds. We therefore rerun the design task with the additional structural constraints. The top three molecules designed by GraphXForm in these runs are shown in the top row of Figure 5. With constraints, we achieve a ‘mean best’ value of $8.23\pm 0.00$ and a ‘mean top 20’ value $7.98\pm 0.02$, which are only slightly lower than the values obtained without structural constraints. We also repeated these constrained design runs with REINVENT-Transformer, given its conceptual similarity to GraphXForm. However, REINVENT-Transformer only reached a ‘mean best’ value of $7.03\pm 1.14$, with its overall best molecule having an objective value of 7.87.
The top three molecules identified by GraphXForm with the structural constraints for the TMB/DMBA task are shown in the bottom row of Figure 5. Under these constraints, GraphXForm achieves a ‘mean best’ value of $8.65\pm 0.00$ and a ‘mean top 20’ value of $8.42\pm 0.03$. Notably, these results remain superior to those of the other methods that do not impose structural constraints listed in Table 1.
This further demonstrates GraphXForm’s flexibility and its ability to consistently design promising molecules.
Figure 5: Top three molecules designed by GraphXForm where: ring size is limited to a maximum of six atoms, any bond between oxygen atoms is disallowed, and single bonds between nitrogen atoms are disallowed.
3.7 Starting molecule design from initial structures
Figure 6: Best molecule designed by GraphXForm when adding a maximum of 3, 5, and 7 atoms to a specific initial structure. We note that increasing the order of a bond within the initial structure is allowed. For TMB/DMBA, we also enforce that the designed molecule stays an alcohol.
One advantage of GraphXForm, which is not feasible when constructing SMILES strings via next-token prediction due to the linear construction, is the ability to start the molecule design process from a predefined structure. This feature is particularly useful when a known candidate, potentially in practical use, possesses desirable properties but could benefit from slight modifications to improve the objective function. To demonstrate this, we consider the IBA task and begin with the best molecule from COSMObase 2020, which has an objective value of 5.57 (see Figure 4). We conduct three individual experiment runs, allowing GraphXForm to add up to 3, 5, and 7 atoms to the initial structure (and alter bond orders as needed). We also maintain the structural constraints outlined in Section 3.6. The original molecule, along with the best molecule from each of the three runs, is shown in the first row of Figure 6. In all cases, the designed molecules exhibit improvements in the objective value.
We repeat a similar experiment for the TMB/DMBA task. During the list screening, we observed that the top three structures were all long-chain alcohols. However, long-chain alcohols tend to have melting points above room temperature, making them unsuitable as solvents for the processes under consideration. The melting point can be lowered by adding branches and increasing bond orders. Thus, we initiate three runs starting from the best molecule identified in the list screening, allowing the addition of up to 3, 5, or 7 atoms, respectively. Additionally, we constrain the design process to ensure the resulting molecule remains an alcohol by preventing any actions from being taken on the hydroxy group (one oxygen bonded to one hydrogen). The results are shown in the second row of Figure 6. While no better molecule is found when adding only up to 3 atoms, a notable improvement occurs when allowing to add 5 or 7 atoms.
This approach provides a powerful tool for situations where it is preferable not to start the design process from scratch but to make targeted modifications to existing molecules.
We note that allowing the agent to remove atoms or bonds would not provide any benefit when starting the design from scratch, i.e., from a single atom. In fact, it would introduce additional redundancies into the search space. However, in the discussed scenarios where the design begins with an existing molecule, enabling the agent to remove atoms or bonds could in fact offer an advantage, as it allows for greater flexibility and deviations from the initial structure. We leave the exploration of this possibility for future work.
Conclusion
We presented GraphXForm, a method for molecular design that follows the successful paradigm of self-supervised pretraining followed by (RL-based) fine-tuning, but operates directly on molecular graphs. By doing so, we addressed challenges faced by string-based methods, such as chemical validity or accomodating structural constraints. We introduced a technique derived from self-improvement learning to stably fine-tune a deep graph-transformer model with eight layers. On two solvent design tasks, we showed that GraphXForm can outperform state-of-the-art molecule design techniques. Additionally, our approach can flexibly adhere to specified structural constraints, such as bond types and functional groups, and can adaptively start the design process from existing molecular structures.
Looking ahead, several promising avenues for future development exist. First, we could expand GraphXForm to include more atoms and ionization states, as well as pretrain the network on much larger databases. We note that these features can already be integrated within our current framework; however, we restricted the set of possible atoms to ensure comparibility between our and other methods. We could further evaluate GraphXForm on additional solvent design tasks, as well as more extensively researched drug design tasks.
We also aim to incorporate more constraints to ensure the suitability of solvents for liquid-liquid separation processes, recognizing that numerous factors – such as boiling and melting points – are critical to a solvent’s effectiveness. This, however, will depend on the availability of reliable property predictors.
Finally, since many structural constraints (e.g., presence of certain atoms, bonds, or formal groups) on the molecular graph can be flexibly formulated and implemented in a general manner within the current framework, we envision integrating GraphXForm with large language models to create a user-friendly design interface. This would allow researchers to formulate constraints in natural language, which would then be translated into the appropriate configuration for the model.
Author contributions
J.P.: Conceptualization; Methodology - Concept and implementation of environment, network architecture, learning algorithm, adaptation of benchmark methods; Analysis; Validation; Visualization; Writing - Original Draft. J.G.R.: Conceptualization; Methodology - Concept and implementation of objective functions and surrogate models, adaptation of benchmark methods; Analysis; Validation; Writing - Original Draft. A.B.W.: Conceptualization; Methodology - Concept and implementation of environment and objective functions, adaptation of benchmark methods; Analysis; Validation; Visualization; Writing - Original Draft. M.G.: Resources; Funding acquisition; Writing - Review & Editing. J.B.: Conceptualization; Resources; Supervision; Project administration; Funding acquisition; Writing - Review & Editing. A.M.: Conceptualization; Resources; Supervision; Project administration; Funding acquisition; Writing - Review & Editing. D.G.G.: Conceptualization; Resources; Supervision; Project administration; Funding acquisition; Writing - Review & Editing.
Data availability
All our code and data for pretraining and fine-tuning GraphXForm is available at https://github.com/grimmlab/graphxform.
Acknowledgements
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 466417970 and 466387255 – within the Priority Programme “SPP 2331: Machine Learning in Chemical Engineering”. This work was performed as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE). Simulations were performed with computing resources granted by RWTH Aachen University under project “thes1232”. The authors gratefully acknowledge the Competence Center for Digital Agriculture (KoDA) at the University of Applied Sciences Weihenstephan-Triesdorf for providing additional computational resources.

### v2

@@
-GraphXForm: Graph transformer for computer-aided molecular design with application to extraction
+GraphXForm: Graph transformer for computer-aided molecular design
 Abstract
-Generative deep learning has become pivotal in molecular design for drug discovery and materials science.
+Generative deep learning has become pivotal in molecular design for drug discovery, materials science, and chemical engineering.
 A widely used paradigm is to pretrain neural networks on string representations of molecules and fine-tune them using reinforcement learning on specific objectives.
@@
 Our approach iteratively modifies a molecular graph by adding atoms and bonds, which ensures chemical validity and facilitates the incorporation of structural constraints.
-We present GraphXForm, a decoder-only graph transformer architecture, which is pretrained on existing compounds and then fine-tuned using a new training algorithm that combines elements of the deep cross-entropy method with self-improvement learning from language modeling, allowing stable fine-tuning of deep transformers with many layers.
-We evaluate GraphXForm on two solvent design tasks for liquid-liquid extraction, showing that it outperforms four state-of-the-art molecular design techniques, while it can flexibly enforce structural constraints or initiate the design from existing molecular structures.
-Machine Learning, ICML
+We present GraphXForm, a decoder-only graph transformer architecture, which is pretrained on existing compounds and then fine-tuned using a new training algorithm that combines elements of the deep cross-entropy method and self-improvement learning.
+We evaluate GraphXForm on various drug design tasks, demonstrating superior objective scores compared to state-of-the-art molecular design approaches.
+Furthermore, we apply GraphXForm to two solvent design tasks for liquid–liquid extraction, again outperforming alternative methods while flexibly enforcing structural constraints or initiating design from existing molecular structures.
+Machine Learning, Molecule Design
 […]
@@
 To address the challenge of capturing long-range dependencies in sequential data that RNNs face, transformers (Vaswani, 2017) have been successfully applied due to their ability to model long-range dependencies more efficiently and as they are widely used in language modeling today (Xu et al., 2024).
-However, transformers are resource-intensive, and needing to fine-tune them with RL can limit the size that is used in practice (Xu et al., 2024; Henderson et al., 2018; Schulman et al., 2017).
+However, transformers are resource-intensive, and fine-tuning them with RL further constrains the model sizes that are practical in real-world applications (Xu et al., 2024; Henderson et al., 2018; Schulman et al., 2017).
 In general, chemical language models may propose string representations of molecules with invalid chemical structures – for example, when SMILES syntax or valence constraints are violated – which has led to numerous works aimed at circumventing this problem (O’Boyle & Dalke, 2018; Krenn et al., 2020; Cheng et al., 2023; Dai et al., 2018).
 Despite recent evidence that invalid SMILES can actually be beneficial from a language modeling perspective (Skinnider, 2024), they can harm the RL component of the pipeline as they can increase sample complexity and necessitate reward shaping to account for them.
-Additionally, the sequential nature of string synthesis makes it difficult to enforce structural constraints, such as ensuring the presence of a minimum number of specific atom types, restrictions on bonds, or even enforcing the presence of certain substructures within the molecules.
+Moreover, the sequential nature of string synthesis makes it challenging to enforce structural constraints – such as ensuring a minimum number of specific atom types, restricting bonding patterns, or initiating design from predefined molecular substructures – often requiring scaffold-constrained techniques (Langevin et al., 2020; Arús-Pous et al., 2020).
 An alternative approach is to represent molecules directly as graphs, where atoms are nodes and bonds are edges, and to develop models that modify a molecular graph or design it directly (Wang et al., 2023; Li et al., 2018; Jin et al., 2018; Zhou et al., 2019; Jensen, 2019; De Cao & Kipf, 2018; Zang & Wang, 2020; Zhang et al., 2023b; Mahmood et al., 2021; Maziarka et al., 2019; Rittig et al., 2023d).
-Working at the graph level, allows explicit encoding of atomic interactions and bonding rules, ensuring chemical validity, and makes it straightforward to start from existing structures and modify them.
+Working at the graph level allows explicit encoding of atomic interactions and bonding rules, ensuring chemical validity, and makes it straightforward to start from existing structures and modify them.
 For example, graph-based methods like GraphGA (Jensen, 2019) employ genetic algorithms to modify molecular graphs directly, even outperforming several neural-based string-synthesis methods without relying on neural networks.
@@
 We propose to combine and extend the strengths of both paradigms: employing transformers for their ability to capture long-range dependencies in sequence data, and leveraging pretraining on existing molecules, all while working directly on the molecular graph.
-More specifically, we use graph transformers (Ying et al., 2021; Maziarka et al., 2019) and formulate molecule design as a sequential task, where a molecular graph – starting from an arbitrary structure – is iteratively extended by placing atoms, bonding them to existing atoms, and modifying the graph by increasing bond orders.
-To this end, we introduce GraphXForm, a decoder-only graph transformer architecture that guides these incremental decisions, predicting the next action based on the current molecular graph.
-Pretrained on existing compounds, GraphXForm employs a novel fine-tuning approach that combines elements of the deep cross-entropy method (Wagner, 2021) and self-improvement learning (Huang et al., 2023; Pirnay & Grimm, 2024a).
-Unlike commonly used deep reinforcement learning methods like REINFORCE (Williams, 1992), this approach allows for stable training of deep transformers with many layers.
-We test GraphXForm for the design of solvents, which are highly relevant for chemical processes.
-While algorithmic advances in molecular design have primarily focused on de novo drug development, other areas such as materials science and chemical engineering have only recently gained attention.
-Recent applications include catalyst (Schilter et al., 2023), fuel (Rittig et al., 2023d; Sarathy & Eraqi, 2024), polymer (Jiang et al., 2024; Yue et al., 2024), surfactant (Nnadili et al., 2024), chemical reaction substrate (Nigam et al., 2023), and solvent (König-Mattern et al., 2024) design.
-We focus specifically on solvents due to their vital role in industrial chemical processes such as reactions, separations, and extractions.
+More specifically, we use graph transformers (Ying et al., 2021; Maziarka et al., 2019) and formulate molecule design as a sequential task, where a molecular graph – starting from an arbitrary structure – is iteratively extended by placing atoms and adding bonds.
+To this end, we introduce GraphXForm, a decoder-only graph transformer architecture that guides these incremental decisions, predicting the next modification based on the current molecular graph.
+Pretrained on existing compounds, we propose a fine-tuning approach for downstream tasks that combines elements of the deep cross-entropy method (Wagner, 2021) and self-improvement learning (Pirnay & Grimm, 2024b).
+Unlike commonly used deep RL methods like REINFORCE (Williams, 1992), this approach allows for stable training of deep transformers with many layers.
+We test GraphXForm for two molecular applications: (1) drug development and (2) solvent design.
+While algorithmic advances in molecular design have been primarily focused on de novo drug development and corresponding benchmarks (Brown et al., 2019; Polykovskiy et al., 2020; Gao et al., 2022; Thomas et al., 2024), other areas such as materials science and chemical engineering have recently gained attention.
+Recent applications include catalyst (Schilter et al., 2023), fuel (Rittig et al., 2023d; Sarathy & Eraqi, 2024), polymer (Jiang et al., 2024; Yue et al., 2025), surfactant (Nnadili et al., 2024), chemical reaction substrate (Nigam et al., 2023), and solvent (König-Mattern et al., 2024) design.
+Thus, we test GraphXForm in both established and newer application areas.
+First, we consider drug design by evaluating GraphXForm on the goal-directed tasks of the well-established GuacaMol benchmark (Brown et al., 2019).
+The benchmark includes multiple design tasks such as drug rediscovery, isomer identification, and multi-property optimization.
+Based on these different design objectives, we demonstrate the soundness of GraphXForm and its competitive performance with state-of-the-art molecular design approaches.
+Secondly, we apply GraphXForm for the design of solvents, which play a vital role in industrial chemical processes such as reactions, separations, and extractions.
 While König-Mattern (König-Mattern et al., 2024) recently applied a graph-based genetic algorithm for solvent design, the use of generative ML-based approaches remains underexplored.
-Therefore, our goal is to compare newer design methods and expand their capabilities.
-We evaluate GraphXForm on two liquid-liquid extraction tasks.
+Therefore, our goal is to compare newer design methods and expand their capabilities by considering molecular structure constraints.
+Specifically, we evaluate GraphXForm on two liquid-liquid extraction tasks.
 The objective function in these tasks is defined by a separation factor based on activity coefficients at infinite dilution.
-We compare GraphXForm on these two downstream tasks to four state-of-the-art molecular design approaches (Graph GA (Jensen, 2019), REINVENT-Transformer (Xu et al., 2024), Junction Tree VAE (Jin et al., 2018), and STONED (Nigam et al., 2021)).
-Additionally, we demonstrate GraphXForm’s unique flexibility and stability by incorporating structural constraints conceptually suited for solvent design, such as preventing certain bond types or preserving complex molecular substructures, allowing us to design solvents with highly tailored properties.
-This capability highlights the strength of our approach in tackling design tasks that are difficult or infeasible for existing methods.
+For both downstream tasks, we compare GraphXForm to state-of-the-art molecular design approaches (Graph GA (Jensen, 2019), REINVENT-Transformer (Xu et al., 2024), Junction Tree VAE (Jin et al., 2018), and STONED (Nigam et al., 2021)).
+Additionally, we demonstrate GraphXForm’s flexibility and stability by incorporating structural constraints conceptually suited for solvent design, such as preventing certain bonding patterns or preserving molecular substructures, allowing GraphXForm to propose design candidates with highly tailored properties.
+This capability highlights the strength of our approach in tackling design tasks that are difficult for existing methods.
 Our contributions are summarized as follows:
   • We formulate molecular design as a sequential task, where an initial structure (e.g., a single atom) is iteratively modified by adding atoms and bonds.
-  • We introduce a graph transformer architecture that takes a molecular graph as input and outputs probability distributions for atom placement and bond modification.
+  • We introduce a graph transformer architecture that takes a molecular graph as input and outputs probability distributions for atom and bond placement.
 This approach maintains the notion of using transformers for molecular design while moving away from string-based methods.
   • We propose a training algorithm that enables the stable and efficient fine-tuning of deep graph transformers on downstream tasks.
-  • We demonstrate that our method outperforms state-of-the-art molecular design techniques on two solvent design tasks and can be easily adapted to satisfy structural constraints or preserve specific molecular structures.
+  • We demonstrate that our method outperforms state-of-the-art molecular design techniques on well-established drug design benchmarks and two solvent design tasks.
+  • We show that our method can be easily adapted to meet structural constraints by preserving or excluding specific molecular moieties and starting the design from initial structures.
 Our code for pretraining and fine-tuning is available at:
@@
 We denote by $\mathcal{M^{\prime}}$ the space of all molecular graphs $m=(\bm{a},\bm{B})$ as described above.
-Accordingly, let $\mathcal{M}\subset\mathcal{M^{\prime}}$ be the subspace of molecular graphs that are chemically valid, i.e., all non-hydrogen atoms follow the octet rule.
+Accordingly, let $\mathcal{M}\subset\mathcal{M^{\prime}}$ be the subspace of molecular graphs that are chemically valid, that is, all non-hydrogen atoms follow the octet rule.
 We refer to following the octet rule as satisfying the ’valence constraints’ throughout the paper.
+In this framework, ionization can be incorporated straightforwardly by extending the alphabet with additional atom types that allow for an adjusted number of non-hydrogen bonds.
+For instance, to model ionized carbon atoms, we can include the types $\mathrm{C^{+}}$ and $\mathrm{C^{-}}$ in $\Sigma$, which are designated with a positive or negative charge and are permitted to form up to five and three non-hydrogen bonds, respectively.
+Chirality can be handled in an analogous manner.
+We refer to the Supplementary Information for the complete alphabet used.
 Sequential molecular graph design
 Similar to Zhou et al.
-(2019), we pose molecular design as a sequential Markov decision process (MDP), where an agent assembles a molecular graph by iteratively adding atoms or increasing the bond order between atoms.
-We note that in the graph, hydrogens are always only considered implicitly, and an addition of an atom or a bond leads to a replacement of an implicit hydrogen.
-On a high level, a single molecule is constructed by the agent as follows: The agent observes an initial molecule $m^{(0)}$, and then performs some action $x^{(0)}$ on it, which results in a new molecule $m^{(1)}$.
+(2019), we pose the molecular design as a sequential Markov decision process (MDP), where an agent assembles a molecular graph by iteratively adding atoms or bonds between atoms.
+We note that in the graph, hydrogens are always only considered implicitly, and an addition of an atom or a bond leads to a replacement of implicit hydrogen.
+On a high level, a single molecule is constructed by the agent as follows: The agent observes an initial molecule $m^{(0)}$ and then performs some action $x^{(0)}$ on it, resulting in a new molecule $m^{(1)}$.
 Then, the agent observes $m^{(1)}$, decides on an action $x^{(1)}$ leading to molecule $m^{(2)}$, and so on.
-This iterative design process terminates at some molecule $m^{(T)}$ once the agent decides to perform a special DontChange action, which does not alter the molecular graph but rather marks the design as finished.
-We formalize this as follows: Let $m^{(t)}=(\bm{a}^{(t)},\bm{B}^{(t)})$ be some molecule with atoms $\bm{a}^{(t)}\in\{1,\dots,k\}^{n^{(t)}}$ and bond matrix $\bm{B}^{(t)}\in\mathbb{N}_{0}^{n^{(t)}\times n^{(t)}}$ as above.
+This iterative design process ends on some molecule $m^{(T)}$ once the agent decides to perform a special action DontChange, which does not alter the molecular graph but rather marks the design as completed.
+We formalize this as follows.
+Let $m^{(t)}=(\bm{a}^{(t)},\bm{B}^{(t)})$ be some molecule with atoms $\bm{a}^{(t)}\in\{1,\dots,k\}^{n^{(t)}}$ and bond matrix $\bm{B}^{(t)}\in\mathbb{N}_{0}^{n^{(t)}\times n^{(t)}}$ as above.
 We transition to a new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$ by performing some action $x^{(t)}\in\mathcal{A}$ on $m^{(t)}$.
@@
   2. 2.
-    $x^{(t)}=\textsc{AddAtom}(j,l)$, with $j\in\{1,\dots,k\}$ and $l\in\{1,\dots,n^{(t)}\}$.
-This action adds an atom of type $\Sigma_{j}$ to the graph and bonds it to the $l$-th atom.
+    $x^{(t)}=\textsc{AddAtom}(j,l,o)$, with $j\in\{1,\dots,k\}$, $l\in\{1,\dots,n^{(t)}\}$, and $o\in\mathbb{N}$.
+This action adds an atom of type $\Sigma_{j}$ to the graph and connects it to the $l$-th atom with a bond of order $o$.
 In particular, we set $n^{(t+1)}=n^{(t)}+1$.
 For the new molecule $m^{(t+1)}=(\bm{a}^{(t+1)},\bm{B}^{(t+1)})$, the atom vector $\bm{a}^{(t+1)}\in\mathbb{N}^{n^{(t+1)}}$ is obtained by appending $j$ to $\bm{a}^{(t)}$.
-The second entry $l\in\{1,\dots,n^{(t)}\}$ indicates that we bond this new $(n+1)$-th atom to the $l$-th atom, i.e., $\bm{B}^{(t+1)}\in\mathbb{N}_{0}^{n^{(t+1)}\times n^{(t+1)}}$ is obtained by appending a zero row and column to $\bm{B}^{(t)}$ and setting $B^{(t+1)}_{n+1,l}=B^{(t+1)}_{l,n+1}=1$.
+The second entry $l\in\{1,\dots,n^{(t)}\}$ indicates that we bond this new $(n+1)$-th atom to the $l$-th atom with order $o$, i.e., $\bm{B}^{(t+1)}\in\mathbb{N}_{0}^{n^{(t+1)}\times n^{(t+1)}}$ is obtained by appending a zero row and column to $\bm{B}^{(t)}$ and setting $B^{(t+1)}_{n+1,l}=B^{(t+1)}_{l,n+1}=o$.
   3. 3.
-    $x^{(t)}=\textsc{IncBond}(j,l)$, meaning that we are increasing the bond order between atoms $j$ and $l$.
+    $x^{(t)}=\textsc{AddBond}(j,l,o)$, meaning that we are adding a bond of order $o\in\mathbb{N}$ between existing, unbonded atoms $j$ and $l$.
 In particular, we have $n^{(t+1)}=n^{(t)}$ and $\bm{a}^{(t+1)}=\bm{a}^{(t)}$.
-The bond matrix $\bm{B}^{(t+1)}$ is obtained from $\bm{B}^{(t)}$ by setting $B^{(t+1)}_{j,l}=B^{(t+1)}_{l,j}=B^{(t)}_{j,l}+1$.
-Note that this allows not only higher order bonds, but also bonds between previously unbonded atoms.
+The bond matrix $\bm{B}^{(t+1)}$ is obtained from $\bm{B}^{(t)}$ by setting $B^{(t+1)}_{j,l}=B^{(t+1)}_{l,j}=o$.
 Given a molecule $m^{(0)}$ and a sequence of actions $x^{(0)},\dots,x^{(t-1)}$, we will also write $(m^{(0)},x^{(0)},\dots,x^{(t-1)})$ for the molecule $m^{(t)}\in\mathcal{M}$ that results from $m^{(0)}$ by sequentially applying the actions $x^{(0)},\dots,x^{(t-1)}$ to $m^{(0)}$.
-We note that, in general, the action sequence $x^{(0)},\dots,x^{(t-1)}$ is not unique to from $m^{(0)}$ to $m^{(t)}$.
-Starting from a molecule in $\mathcal{M}$, by removing chemically invalid actions from the action space (which would lead to violation of valence constraints) at each step, we can guarantee to stay in the space of chemically valid molecules.
-Once the agent chooses the DontChange action, the design process is considered finished.
-We note that, starting from an appropriate initial atom, it is in fact possible to reach every target molecule in the chemically valid space $\mathcal{M}$ (i.e., all molecular graphs that can be constructed from the alphabet $\Sigma$ and that satisfy the valence constraints) when starting from any atom that exists in the target molecule.
+In general, the action sequence $x^{(0)},\dots,x^{(t-1)}$ is not unique to get from $m^{(0)}$ to $m^{(t)}$.
+Starting from a molecule in $\mathcal{M}$, by removing chemically invalid actions from the action space (which would lead to violation of valence constraints) at each step, we can guarantee staying in the space of chemically valid molecules.
+Once the agent chooses the action DontChange, the design process is considered complete.
+We note that starting from an appropriate initial atom, it is in fact possible to reach every target molecule in the chemically valid space $\mathcal{M}$ (i.e., all molecular graphs that can be constructed from the alphabet $\Sigma$ and that satisfy the valence constraints) when starting from any atom that exists in the target molecule.
 For an illustrative molecule construction, see Fig.
 1.
-Each addition of a bond or atom removes an implicit hydrogen from the respective atom, allowing for a straightforward representation of ionization states: Ions can be incorporated simply by adding charged atom types to the predefined alphabet, without the need for additional modifications to the action space.
-In contrast, the graph-based approach can not handle stereochemistry in a straightforward way.
-Figure 1: Example for the sequential application of actions $x^{(0)},\dots,x^{(3)}$ to a molecule $m^{(0)}$, using the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$.
+Figure 1: Example for the sequential application of actions $x^{(0)},x^{(1)},x^{(2)}$ to a molecule $m^{(0)}$, using the alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$.
 We show the index of each atom, which can be arbitrarily chosen at the beginning.
 Light blue indicates where in the graph an action is applied.
-The last action is DontChange, which does not change the molecular graph, but marks it as a finished design.
+The last action is DontChange, which does not change the molecular graph, but marks it as a complete design.
 Molecular optimization
@@
 That means, that we set the corresponding probability to zero in the distribution predicted by the policy (before re-normalizing the distribution).
-Our graph-based approach allows us to extend this concept of action masking to enforce additional constraints, such as particular bonding patterns, minimum/maximum number of atoms and their types, or the presence of structural motifs like rings.
+Our graph-based approach allows us to extend this concept of action masking to enforce additional constraints, such as particular bonding patterns, minimum/maximum number of atoms and their types, or restricting structural motifs like rings.
 We explore these constraints in detail in Section 3.
 We note that it is possible to simply assign an objective value of $\infty$- ∞ to molecules that violate these constraints after the design process.
-However, the ability to preemptively avoid constraint-violating regions by masking actions during the design process can significantly reduce the search space.
-In contrast, methods that generate molecules by sequentially synthesizing SMILES (Gupta et al., 2018; Xu et al., 2024; Mazuz et al., 2023) strings struggle to enforce such constraints during the construction process while also adhering to the grammatical rules of SMILES.
+However, the ability to preemptively avoid constraint-violating regions by masking actions during the design process reduces the search space.
 2.2 Learning algorithm
 We now introduce our method for training the policy neural network $\pi_{\theta}$.
-Our proposed learning algorithm is a hybrid of the deep cross-entropy method (CEM) (Wagner, 2021) and the self-improvement learning (SIL) (Huang et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024a, b) strategy.
-Both approaches train the network over multiple epochs in a self-improving loop, where, in each epoch, the current policy is used to generate a set of action sequences.
-The best sequences are selected as ’pseudo-labels’, i.e., they serve as the dataset for supervised training.
-The network is then trained for one epoch to assign higher probabilities to these sequences before the process is repeated with the updated network to generate new action sequences.
-Unlike other deep RL algorithms, this approach does not require reward shaping or value approximation.
-Furthermore, training in a supervised way provides stability and allows for the use of larger, decoder-only transformer architectures (Vaswani, 2017), which would be prohibitively slow under standard RL.
-However, SIL and the deep CEM differ in key aspects: The deep CEM is formulated for problems with a single instance (as in our molecular design task) and retains a certain percentage of the best action sequences in each epoch, which are obtained through simple sampling from the policy’s predicted distributions at each step.
-SIL, on the other hand, is typically applied to problems with infinitely many instances and employs more advanced sequence decoding techniques, such as sequence sampling without replacement (Kool et al., 2019), to improve solution quality and diversity while maintaining efficient decoding speed.
-In our proposed approach, we aim to develop a method that works for single-instance problems, as in the deep CEM, while retaining the diverse sampling mechanism used in SIL methods (Pirnay & Grimm, 2024a, b).
-The pseudocode for our approach is presented in Algorithm 1, and we describe the key steps below:
+Our proposed learning algorithm is a hybrid of the deep cross-entropy method (CEM) (Wagner, 2021) and self-improvement learning (SIL) (Huang et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024a, b), a sampling-based approach to expert iteration (Anthony et al., 2017).
+Both the deep CEM and SIL train the neural network over multiple epochs in a self-improving loop.
+In each epoch, the current policy is used to generate a set of action sequences, from which the best sequences are selected as ’pseudo-labels’ to serve as the dataset for supervised training.
+The network is then trained for one epoch to assign higher probabilities to these sequences, and the updated network is subsequently used to generate new action sequences.
+Unlike many deep RL algorithms, this approach does not require reward shaping or value approximation.
+Moreover, training in a supervised way provides stability and facilitates the use of larger, decoder-only transformer architectures (Vaswani, 2017).
+There are key differences between SIL and the deep CEM.
+The deep CEM is formulated for problems with a single instance (as in our molecular design task) and retains a fixed percentage of the best action sequences in each epoch, which are obtained through simple sampling from the policy’s predicted distributions at each step.
+In contrast, SIL is typically applied to problems with infinitely many instances and employs more advanced sequence decoding techniques – such as sequence sampling without replacement (Kool et al., 2019) – to improve solution quality and diversity while maintaining efficient decoding speed.
+In our proposed approach, we seek to develop a method that works for single-instance problems, as in the deep CEM, while retaining the diverse sampling mechanism used in SIL methods (Pirnay & Grimm, 2024a, b).
+In particular, we adopt the Take a Step and Reconsider (TASAR) method (Pirnay & Grimm, 2024b) to sample action sequences, as detailed below.
+The pseudocode for our approach is presented in Algorithm 1, and we describe the key steps as follows:
   1. 1.
@@
   2. 2.
-    Throughout the training process, we maintain a set BestFound of the best molecules discovered, which will be returned at the end. Initially, this set contains only the starting molecule $m_{0}$.
+    Throughout training, we maintain a set BestFound containing the best molecules discovered so far. This set is initially initialized with the starting molecule $m_{0}$.
   3. 3.
-    In each epoch, we incrementally sample action sequences without replacement from the policy until either an improved molecule is found or a maximum number of samples is reached (lines 5-10).
-For sampling sequences without replacement, we use the efficient batch-wise incremental method introduced by Shi et al.
-(2020), which relates to Stochastic Beam Search (Kool et al., 2019) and is employed in SIL methods (Pirnay & Grimm, 2024a, b).
-Sampling without replacement ensures that unique action sequences are generated, exploring the search space effectively (especially for shorter action sequences (Shi et al., 2020)).
-We emphasize that, while different action sequences may lead to the same molecule, this does not pose an issue because the policy is only concerned with generating the action sequences rather than the molecular structure.
-Since the goal of sampling sequences without replacement is to produce distinct sequences, the policy still generates diverse outputs, even if the resulting molecules can be identical.
+    In each epoch, we sample action sequences from the policy using the TASAR method (Pirnay & Grimm, 2024b).
+Specifically, $\beta$ action sequences are sampled without replacement using Stochastic Beam Search (Kool et al., 2019) with a beam width $\beta\in\mathbb{N}$.
+These sequences are then evaluated using the objective function.
+The best action sequence among the $\beta$ samples is followed for a predefined number $\sigma$ of actions.
+Subsequently, alternative, previously unseen actions are sampled from the resulting partial sequence to further explore and potentially improve the solution, and the process continues.
+Sampling without replacement ensures that unique action sequences are generated, effectively exploring the search space (particularly for shorter sequences (Shi et al., 2020)).
+Although different action sequences may result in the same molecule, this does not pose an issue for TASAR since the policy is concerned with generating action sequences rather than the molecular structure itself.
+By sampling sequences without replacement, the policy is encouraged to produce diverse outputs, even if the resulting molecules are identical.
   4. 4.
-    After sampling, the set BestFound serves as a training dataset (lines 12-14).
-Similar to how decoder-only models in language modeling are trained to predict the next token from partial text, we sample batches of intermediate molecules and train the network with a cross-entropy loss to predict the next action for the corresponding molecule.
+    After sampling, the set BestFound serves as the training dataset for supervised learning (lines 6-8).
+Similarly to how decoder-only models in language modeling are trained to predict the next token from partial text, we sample batches of intermediate molecules and train the network with a cross-entropy loss to predict the next action for the corresponding molecule.
   5. 5.
@@
 Input: $s\in\mathbb{N}$: number of best molecules to keep
-Input: $q_{\min},q_{\max}\in\mathbb{N}$: minimum and maximum number of molecules to sample without replacement
+Input: $\beta,\sigma\in\mathbb{N}$: beam width and step size (hyperparams for TASAR (Pirnay & Grimm, 2024b))
 1 $\textsc{BestFound}\leftarrow\{m_{0}\}$
@@
 3 foreach epoch do
-4       $\textsc{Sampled}\leftarrow\emptyset$
-5       while $|\textsc{Sampled}|<q_{\max}$ do
-6             $m\leftarrow$ sample from $\mathcal{M}$ with probability $\pi_{\theta}\left(m\ |\ m_{0},m\notin\textsc{Sampled}\right)$
-7             $\textsc{Sampled}\leftarrow\textsc{Sampled}\cup\{m\}$
-8             if $f(m)>\textsc{BestObj}$ then
-9                   $\textsc{BestObj}\leftarrow f(m)$
-10                   if $|\textsc{Sampled}|>q_{\min}$ then
-11                         break
-12
-13
+[…]
 […]
@@
 8             minimize batch-wise cross-entropy loss $L_{\theta}=-\frac{1}{B}\sum_{j=1}^{B}\log\pi_{\theta}\left(x_{t_{j}}^{(j)}\ |\ m_{t_{j}}^{(j)}\right)$
-18
-19
+9
+10
 return BestFound
@@
 2.3 Policy network architecture
-The policy network takes a molecule as input and predicts a probability distribution over next actions.
-To achieve this, we use a simplified version of the Graphormer (Ying et al., 2021) architecture, which treats the molecule’s atoms as an unordered sequence of nodes.
-This sequence is passed through a stack of transformer layers, with the attention mechanism augmented by bonding information.
+The policy network receives a molecule as input and predicts a probability distribution over possible next actions using a simplified version of the Graphormer (Ying et al., 2021).
+This model treats the molecule’s atoms as an unordered set of nodes and processes them through a stack of transformer layers with the attention mechanism augmented by bonding information.
 The resulting latent representations of the nodes are then used to predict action distributions.
+Splitting actions
+To make the network’s predictions more fine-grained, we decompose each action (except DontChange into three sub-actions, each determined by a separate forward pass through the network:
+  • Action level 0: The agent decides whether to end the design process (choosing DontChange) or to modify the molecule.
+In the latter case, it selects the first atom $j$, which can be an already present one for $\textsc{AddBond}(j,l,o)$ or a new one from the alphabet for $\textsc{AddAtom}(j,l,o)$.
+  • Action level 1: Given a modification is intended, the agent selects a second atom $l$ from the current molecule, indicating a new bond between $j$ and $l$.
+  • Action level 2: Finally, the agent determines the order $o$ of the bond between $j$ and $l$.
 Figure 2 provides an illustration, and we elaborate on the architecture below:
@@
 The molecule’s underlying graph is augmented with a virtual node (indexed with 0) and embedded into the latent space $\mathbb{R}^{d}$.
+Learnable vectors are added to these embeddings to encode the current action level and decisions on previous levels.
 b.
@@
 This virtual bond is represented in the bond matrix by an integer outside the range of standard bond orders.
-The virtual atom functions similarly to the special [CLS]-token in BERT (Devlin, 2018) and accumulates sequence-level information for downstream tasks.
+The virtual atom functions similarly to the special [CLS]-token in BERT (Devlin et al., 2019) and accumulates sequence-level information for downstream tasks.
 Conceptually, the virtual atom acts as an additional message proxy between nodes in the molecular graph.
-The sequence $(a_{0},a_{1},\dots,a_{n})$ is embedded into a corresponding sequence of latent node representations $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1},\dots,\bm{\hat{a}}_{n})\in\mathbb{R}^{d}$, where each $\bm{\hat{a}}_{i}$ is a learnable vector associated with the atom type.
-Additionally, let $\bm{z}_{i}\in\mathbb{R}^{d}$ be a learnable embedding that depends on the number of bonds formed by the $i$-th atom.
-The augmented sequence $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1}+\bm{z}_{1},\dots,\bm{\hat{a}}_{n}+\bm{z}_{n})$ is then passed through a stack of transformer layers using ReZero normalization (Bachlechner et al., 2021).
-Importantly, to preserve the permutation equivariance of the network (i.e., the order of the atoms does not matter), no positional encoding is applied.
+The atom sequence $(a_{0},a_{1},\dots,a_{n})$ is first embedded latent representations $(\bm{\hat{a}}_{0},\bm{\hat{a}}_{1},\dots,\bm{\hat{a}}_{n})$ in $\mathbb{R}^{d}$, where each $\bm{\hat{a}}_{i}$ is a learnable vector corresponding to the atom type.
+We then augment these representations with additional learnable embeddings that encode the current action level and other relevant information.
+Specifically, we introduce a learnable vector $\bm{r}\in\mathbb{R}^{d}$ to represent the current action level.
+For each atom $i\in{1,\dots,n}$, we further include the embeddings $\bm{w}_{i}^{(0)}$ and $\bm{w}_{i}^{(1)}$, which indicate whether the $i$-th atom has been selected at action levels 0 and 1, respectively.
+Moreover, let $\bm{z}_{i}\in\mathbb{R}^{d}$ be an embedding that reflects the total number of bonds (i.e., the degree) formed by the $i$-th atom.
+The final augmented sequence
+$(\bm{\hat{a}}_{0}+\bm{r},\bm{\hat{a}}_{1}+\bm{z}_{1}+\bm{w}_{1}^{(0)}+\bm{w}_{1}^{(1)},\dots,\bm{\hat{a}}_{n}+\bm{z}_{n}+\bm{w}_{n}^{(0)}+\bm{w}_{n}^{(1)})$ (3)
+is then passed through a stack of transformer layers using ReZero normalization (Bachlechner et al., 2021).
+Importantly, to maintain permutation equivariance — ensuring that the order of the atoms does not affect the outcome — we do not apply any positional encodings.
 To incorporate bonding information within the transformer layers, let $(\bm{A}_{ij})_{0\leq i,j\leq n}\in\mathbb{R}^{(n+1)\times(n+1)}$ be the computed self-attention matrix in a layer (for any attention head) corresponding to the input sequence $(a_{0},a_{1},\dots,a_{n})$.
@@
 We note that $b_{ij}$ for each bond order is learned independently across layers and attention heads.
-Action distribution
-The stack of transformer layers produces a transformed sequence of node embeddings $(\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n})$, where each $\bm{e}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th atom $a_{i}$ in the original sequence.
-This sequence is used to predict three distributions, $P^{(0)},P^{(1)},P^{(2)}$, which together factorize the full distribution over possible next actions.
-Informally, $P^{(0)}$ represents a distribution over whether to terminate the synthesis, add a new atom, or increase the bond order between two existing atoms.
-The distribution $P^{(1)}$ specifies which atom to select, either as the atom to bond a newly added atom to, or as the first atom of a pair for increasing bond order.
-Finally, $P^{(2)}$ is used in the case of bond increase to specify the second atom in the pair.
-Formally, the unnormalized log-probabilities (logits) for these distributions are obtained as follows:
-  • $P^{(0)}$: We compute the logits for $P^{(0)}=(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k},p^{(0)}_{k+1})$ by projecting the vector $\bm{e}_{0}$ using a linear layer $g_{0}\colon\mathbb{R}^{d}\to\mathbb{R}^{k+2}$.
-Here, $p^{(0)}_{0}$ is the probability of choosing the action DontChange.
-For $1\leq j\leq k$, $p_{j}$ represents the probability of adding an atom of type $\Sigma_{k}$ (i.e., $\textsc{AddAtom}(j,\cdot)$ without specifying the bonding atom yet).
-Lastly, $p^{(0)}_{k+1}$ is the probability of increasing the bond between two atoms ($\textsc{IncBond}(\cdot,\cdot)$) without specifying the atom pair.
-  • $P^{(1)}$: To obtain $P^{(1)}=(p^{(1)}_{1},\dots,p^{(1)}_{n})$, we apply a linear layer $g_{1}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each $\bm{e}_{1},\dots,\bm{e}_{n}$.
-Then, $p^{(1)}_{j}$ represents the probability of selecting the $j$-th atom to bond to in $\textsc{AddAtom}(\cdot,\cdot)$, or as the first atom in $\textsc{IncBond}(\cdot,\cdot)$.
-  • $P^{(2)}$: Similarly, for $P^{(2)}=(p^{(2)}_{1},\dots,p^{(2)}_{n})$, we apply another linear layer $g_{2}\colon\mathbb{R}^{d}\to\mathbb{R}$ to compute the logits for the probability $p^{(2)}_{l}$, representing the likelihood of selecting the $l$-th atom as the second atom in IncBond.
-In practice, during inference, we use $P^{(0)}$ and $P^{(1)}$ directly to choose the first atom for both AddAtom and IncBond, and discard $P^{(2)}$.
-If IncBond has been selected from $P^{(0)}$, we reprocess the molecule through the network, marking the atom chosen from $P^{(1)}$ by adding a distinct learnable embedding to its corresponding sequence element.
-In this second pass, we use only $P^{(2)}$ to select the second atom for IncBond.
-This multi-step action prediction allows us to easily mask actions (i.e., setting their probability to zero in the policy) that would lead to invalid molecules.
+Action level distributions
+The stack of transformer layers outputs a sequence of node embeddings $(\bm{e}_{0},\bm{e}_{1},\dots,\bm{e}_{n})$, where each $\bm{e}_{i}\in\mathbb{R}^{d}$ corresponds to the original atom $a_{i}$.
+These embeddings are then used to simultaneously predict three probability distributions, $P^{(0)}$, $P^{(1)}$, and $P^{(2)}$, each corresponding to one of the action levels.
+The agent selects an action based on the distribution associated with the current action level.
+The unnormalized log-probabilities (logits) for these distributions are computed as follows:
+  • Action level 0 ($P^{(0)}$): We denote the logits for $P^{(0)}$ as
+    $(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k},q^{(0)}_{1},\dots,q^{(0)}_{n})\in\mathbb{R}^{k+1+n}.$ (5)
+    The first $k+1$ logits, $(p^{(0)}_{0},p^{(0)}_{1},\dots,p^{(0)}_{k})$, are obtained by projecting the vector $\bm{e}_{0}$ through a linear layer $g_{0}\colon\mathbb{R}^{d}\to\mathbb{R}^{k+1}$.
+Here, $p^{(0)}_{0}$ corresponds to the DontChange action, and for $1\leq j\leq k$, the logit $p_{j}^{(0)}$ corresponds to adding a new atom of type $\Sigma_{j}$ (i.e., serving as the first parameter $j$ in $\textsc{AddAtom}(j,\cdot,\cdot)$.
+The remaining logits, $q^{(0)}_{1},\dots,q^{(0)}_{n}$, are computed by applying a linear layer $h_{0}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each of the embeddings $\bm{e}_{1},\dots,\bm{e}_{n}$.
+For $1\leq j\leq n$, the logit $q^{(0)}_{j}$ corresponds to selecting the first parameter $j$ for the $\textsc{AddBond}(j,\cdot,\cdot)$ action.
+  • Action level 1 ($P^{(1)}$): The logits $(q^{(1)}_{1},\dots,q^{(1)}_{n})$ for this level are obtained by applying a linear layer $h_{1}\colon\mathbb{R}^{d}\to\mathbb{R}$ independently to each of $\bm{e}_{1},\dots,\bm{e}_{n}$.
+Here, $q^{(1)}_{l}$ represents the choice of the second parameter $l$ for either $\textsc{AddAtom}(j,l,\cdot)$ or $\textsc{AddBond}(j,l,\cdot)$.
+  • Action level 2 ($P^{(2)}$): The logits $(p^{(2)}_{1},\dots,p^{(2)}_{y})$, which correspond to the bond order $o$ in either $\textsc{AddAtom}(j,l,o)$ or $\textsc{AddBond}(j,l,o)$, are computed by projecting $\bm{e}_{0}$ using a linear layer $g_{2}\colon\mathbb{R}^{d}\to\mathbb{R}^{y}$, where $y$ is a predefined maximum bond order.
+After the agent selects an action at the current action level, the chosen information is fed back into the network by updating the learnable vectors $\bm{r},\bm{w}_{i}^{(0)}$ and $\bm{w}_{i}^{(1)}$.
+This updated state is then used for the subsequent forward pass when predicting the next action level.
+The multi-step action prediction allows us to easily mask actions (i.e., setting their probability to zero in the policy) that would lead to invalid molecules.
 While checking for actions that would violate constraints adds a small amount of computational overhead, it has a significant benefit: invalid molecules can be immediately disregarded, preventing the network from wasting resources on infeasible designs.
@@
 3 Results and Discussion
-First, we outline our design tasks, including the objective functions and property prediction methods.
-Next, we describe the experimental setup common to all methods and outline the specific setup and hyperparameters for our method, followed by a brief overview of the methods used for comparison.
-We then present preliminary baseline objectives obtained from screening a list of available molecules and discuss our results in relation to those of the other methods.
-3.1 Example solvent design tasks
+We begin by outlining the experimental setup and hyperparameters for our method.
+Our first case study tackles the goal-directed design tasks from the well-established GuacaMol benchmark (Brown et al., 2019).
+In our second case study, we explore two solvent design tasks by detailing the goals, their underlying objective functions, and the property prediction methods used.
+To contextualize our results, we also present preliminary baselines obtained by screening available molecules and compare our findings with those of other approaches.
+3.1 GraphXForm: Pretraining and fine-tuning
+Network hyperparameters
+For all case studies and experiments, we set the latent space dimension to $d=512$ (see Section 2.3).
+The network comprises ten transformer layers with ReZero normalization, each featuring eight attention heads and a feed-forward dimension of 2048.
+Alphabet and pretraining
+We define the atom alphabet $\Sigma$ to include C, N, O, F, P, S, Cl, Br, and I, along with variants for ionization and chirality (see Supplementary Information for the complete alphabet).
+Although the agent is capable of learning without prior knowledge, we pretrain the network on known molecules using SMILES strings from the ChEMBL database (Davies et al., 2015).
+We filtered the database to include only molecules containing atoms from $\Sigma$ and split it into a training set of approximately 1.5 million molecules and a validation dataset of around 70,000 molecules.
+Each SMILES string is converted into an action sequence in our graph formulation; since the conversion is not unique, we select one possible sequence arbitrarily.
+These sequences are used to train the model in a self-supervised manner – predicting the next action given previous actions – as outlined in lines 6–8 of Algorithm 1.
+Training is performed with a dropout rate of 0.1, a batch size of 512, and over a total of 1.5 million batches.
+Petraining the network on existing molecules establishes a prior that captures the characteristics of real molecules, which is crucial for most generative methods such as REINVENT (Olivecrona et al., 2017; Xu et al., 2024) or JT-VAE (Jin et al., 2018).
+Fine-tuning
+Given an objective function $f\colon\mathcal{M}^{\prime}\to\mathbb{R}\cup\{-\infty\}$, we fine-tune the pretrained policy network using the learning algorithm described in Section 2.2.
+We train only the weights of the final linear layers $g_{0},g_{2},h_{0},h_{1}$.
+Unless stated otherwise, we initialize the molecule $m_{0}$ as a single carbon atom.
+We set $s=100$ as the number of top molecules to retain throughout training.
+For the TASAR sampling procedure, we use a beam width of $\beta=512$ and a step size of $\sigma=12$; that is, after every four added atoms or bonds, TASAR seeks better alternative solutions.
+At the end of each epoch, we train the network on 20 batches of size 64, sampled uniformly from the top 100 molecules.
+We intentionally keep the number of training batches per epoch relatively low to prevent premature overfitting, though in most runs increasing the number of batches does not harm performance and may even speed up convergence.
+Runtime and computational budget
+The runtime of a method can vary significantly due to differences in implementation and available hardware.
+A common approach to align computational resources is to limit the number of calls to the objective function (Gao et al., 2022).
+While this is useful for comparing sample efficiency, it offers limited insight into overall efficiency when objective evaluations are inexpensive.
+In our case studies, objective function evaluations are relatively cheap: For the solvent design tasks, for example, these are computed by evaluating the designed molecule with a surrogate neural network that can process batches in parallel.
+Therefore, to provide a practical comparison between our method and competing approaches, each design run is executed until convergence or until a maximum wall-clock time of eight hours is reached.
+We run all experiments with a single NVIDIA H100 GPU with 80 GB of memory.
+3.2 Case study 1: Drug design
 Objectives
-We aim to design solvents for two-phase aqueous-organic systems used in liquid-liquid extraction.
+We evaluate GraphXForm for de novo molecular design using the 20 goal-directed tasks from the GuacaMol benchmark (Brown et al., 2019).
+These tasks span drug rediscovery, similarity-based design, multi-property optimization (MPO), and scaffold hopping.
+We selected GuacaMol as an established benchmark because our primary focus in this study is to demonstrate the overall optimization capability of our method – i.e., achieving molecules with high scores.
+We note, however, that current comprehensive benchmarking also considers factors such as sample efficiency and molecular diversity (Gao et al., 2022; Thomas et al., 2022, 2024).
+Such aspects could be considered in future applications and extensions of GraphXForm, e.g., by further investigating replay buffers (Guo & Schwaller, 2024a, c) and the TASAR parameters.
+Benchmark methods
+For the drug design tasks, we compare GraphXForm against Graph GA (Jensen, 2019) and REINVENT-Transformer (Xu et al., 2024).
+Graph GA (Jensen, 2019) employs a genetic algorithm that directly operates on the molecular graph, mutating atoms and fragments using crossover rules derived from graph matching.
+This non-learning method is highly effective at making fine-grained local modifications, and it has been shown to outperform several SMILES-based learning approaches (Jensen, 2019; Gao et al., 2022; Brown et al., 2019).
+In contrast, REINVENT-Transformer (Xu et al., 2024) designs molecules by synthesizing SMILES strings in an autoregressive manner.
+This method uses a transformer network that is pretrained in a self-supervised fashion on known molecules and then fine-tuned with reinforcement learning via a variant of the REINFORCE algorithm (Williams, 1992).
+We include REINVENT-Transformer in our comparisons because, like GraphXForm, it relies on pretrained transformers and generates molecules autoregressively.
+However, while REINVENT-Transformer constructs molecules token by token from a predefined vocabulary, GraphXForm operates directly on the molecular graph by adding atoms and bonds.
+GuacaMol results
+Table 1 presents the scores of the best molecules found for four representative (Gao et al., 2024) tasks: Ranolazine MPO, Perindopril MPO, Sitagliptin MPO, and Scaffold Hop.
+A complete list of results across all 20 tasks is provided in Supplementary Table 1.
+For Graph GA, we report the results as originally published by Brown et al.
+(2019).
+For REINVENT-Transformer, we conducted experiments using the source code and pretrained model provided by Gao et al.
+(2022).
+GraphXForm outperforms both Graph GA and REINVENT-Transformer in different drug design tasks.
+As shown in Table 1, GraphXForm finds molecules with significantly higher scores for the three MPO cases compared to Graph GA and REINVENT-Transformer.
+For the scaffold hop task, GraphXForm achieves the best possible score of 1 similar to Graph GA.
+Considering the perfomance across all 20 tasks of the GuacaMol benchmark (see Supplementary Table 1), GraphXForm attains a total summed score of 18.227, compared to 17.983 for Graph GA.
+Furthermore, Supplementary Table 1 demonstrates that our method overall outperforms other classic baselines from the original GuacaMol paper, as well as a recent optimization method that utilizes multiple GPT agents for drug design (Hu et al., 2024).
+These results underscore the robust molecular optimization capabilities of GraphXForm.
+Table 1: Performance of different molecular design methods across four tasks from the goal-directed GuacaMol (Brown et al., 2019) benchmark.
+We report the objective function evaluation of the best molecule found.
+Method Ranolazine MPO Perindopril MPO Sitagliptin MPO Scaffold Hop
+Graph GA  (Jensen, 2019) 0.920 0.792 0.891 1.000
+REINVENT-Transformer  (Xu et al., 2024) 0.934 0.679 0.735 0.582
+GraphXForm (ours) 0.944 0.835 0.965 1.000
+3.3 Case study 2: Solvent design
+Objectives
+To further evaluate GraphXForm beyond already-established drug design benchmarks, we propose a solvent design task for two-phase aqueous-organic systems used in liquid-liquid extraction.
 Our focus is on two examples motivated by biotechnology, where products are typically produced in an aqueous solution using microorganisms or enzymes.
@@
 We assume a spatially uniform temperature of 298 K in both examples.
-The first task focuses on the separation of isobutanol (IBA) from water, a common liquid-liquid extraction process.
+The first solvent design task focuses on the separation of isobutanol (IBA) from water, a common liquid-liquid extraction process.
 The chosen solvent should be largely immiscible with water (i.e., low solubility exhibited for both the solvent in water and water in the solvent) and possess high affinity for IBA.
@@
 Hereby, the solvent-independent constant $\gamma_{\text{IBA,W}}^{\infty}$ is omitted.
-Our second task centers on an extraction process presented by Peters et al.
-(2008), who carried out a solvent screening using COSMO-RS as a property predictor.
+Our second solvent design task centers on an extraction process presented by Peters et al.
+(Peters et al., 2008), who carried out a solvent screening using COSMO-RS as a property predictor.
 Here, an enzymatic reaction in aqueous medium converts 3,5-dimethoxy-benzaldehyde (DMBA) molecules to (R)-3,3’,5,5’-tetra-methoxy-benzoin (TMB).
@@
 Following Peters et al.
-(2008), we maximize the ratio $P_{\text{TMB}}^{\infty}/P_{\text{DMBA}}^{\infty}$, while additionally enforcing the miscibility gap constraint from Equation 6 leading to the following scalar objective:
+(Peters et al., 2008), we maximize the ratio $P_{\text{TMB}}^{\infty}/P_{\text{DMBA}}^{\infty}$, while additionally enforcing the miscibility gap constraint from Equation 8 leading to the following scalar objective:
 $\max\frac{\gamma_{\text{TMB,S}}^{\infty}}{\gamma_{\text{DBMA,S}}^{\infty}}+\left(\tanh\left(\gamma_{\text{S,W}}^{\infty}\cdot\gamma_{\text{W,S}}^{\infty}-\exp(4)\right)-1\right)\cdot 10.$ (11)
 Hereby, we again omitted the constants $\gamma_{\text{TMB,W}}^{\infty}$ and $\gamma_{\text{DBMA,W}}^{\infty}$.
+We note that further thermodynamic properties, e.g., boiling and melting points, and sustainability indicators, such as biodegradability and toxicity, are highly relevant for the practical effectiveness of solvents.
+Such properties could be considered as part of the objective function or as additional constraints in future work.
 Property prediction
 To obtain activity coefficients at infinite dilution, we use a state-of-the-art graph neural network (GNN)  (Gilmer et al., 2017; Reiser et al., 2022; Rittig et al., 2023c).
-Specifically, we employ the Gibbs-Helmholtz (GH-) GNN that was developed by Medina et al.
-(2023) for predicting activity coefficients at infinite dilution of binary mixtures at varying temperature.
+Specifically, we employ the Gibbs-Helmholtz (GH-) GNN  (Medina et al., 2023) that was developed by Sanchez Medina et al.
+for predicting activity coefficients at infinite dilution of binary mixtures at varying temperature.
 The GH-GNN takes the molecular graphs of the two molecules within a binary mixture and the temperature as inputs.
@@
 A large data set of experimental activity coefficients at infinite dilution from the DECHEMA Chemistry Data Series  (Gmehling et al., 2008) was used for training.
-For further details on the GH-GNN architecture, we refer to Medina et al.
-(2023).
+For further details on the GH-GNN architecture, we refer to  (Medina et al., 2023).
 We note that many GNN models and other ML models such as transformers have been developed for activity coefficient prediction  (Rittig et al., 2023a; Winter et al., 2022; Damay et al., 2021), also considering the composition-dependency and thermodynamic consistency  (Rittig et al., 2023b; Winter et al., 2023; Rittig & Mitsos, 2024; Specht et al., 2024).
@@
 This model has shown high prediction accuracy, outperforming well-established methods for predicting activity coefficients at infinite dilution such as UNIFAC  (Fredenslund et al., 1975) or COSMO-RS  (Klamt et al., 2010), cf.
-(Medina et al., 2023).
+Medina et al.
+(2023).
 Future work could investigate further additional activity coefficient models or directly predicting partition coefficients with ML  (Zamora et al., 2023; Nevolianis et al., 2024).
-3.2 General computational setup
-To ensure comparability across methods in the design task, we restrict the set of available atoms to carbon, nitrogen, and oxygen (alphabet $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$).
-In the solvent design tasks we consider, molecules are evaluated using objective functions based on a surrogate model for predicting activity coefficients (see Section 3.1).
-To avoid potential exploitation of the surrogate model by constructing large molecules that are beyond the size of typically used solvents and thus likely to be different from the molecules used to train the surrogate model, we constrain the design to molecules of a moderate size, with a maximum of 25 atoms.
-The runtime of a method can vary significantly due to differences in implementation or available hardware.
-One approach to align computational resources is to limit the number of calls to the objective function (Gao et al., 2022).
-While this method is useful for comparing sample efficiency, it provides limited insight into the overall efficiency of a method when objective function evaluations are relatively inexpensive.
-Given the efficiency of the surrogate model, objective function evaluations can be parallelized and run in batches very quickly.
-Therefore, we take a practical approach to align computational resources: each design run of a method is allocated a time budget of eight wall-clock hours on identical hardware, utilizing a single NVIDIA H100 GPU with 80 GB of memory.
-3.3 Our method: GraphXForm
-Network hyperparameters
-For all our experiments, we set the dimension of the latent space $\mathbb{R}^{d}$ to $d=128$ (see Section 2.3).
-The network consists of eight transformer layers with ReZero normalization, each with eight heads in the multi-head attention, and a feed-forward dimension of 512.
-Pretraining
-While it is possible for the agent to learn without any prior knowledge of existing compounds, we pretrain the network on known molecules using SMILES strings from the ChEMBL database (Davies et al., 2015).
-We filtered the database to ensure the molecules contain only atoms within the alphabet $\Sigma$, yielding approximately 370,000 molecules.
-Each SMILES string is then converted into an action sequence in our graph formulation.
-Since this conversion is not unique, we arbitrarily select one possible action sequence.
-The resulting sequences are used to train the model in a self-supervised way, predicting the next action given the previous actions, as outlined in lines 12-14 in Algorithm 1.
-Training is conducted with a batch size of 128, over a total of 2.5 million batches.
-Petraining the network on existing molecules establishes a prior that captures the characteristics of real molecules, which is crucial for most generative methods such as REINVENT (Olivecrona et al., 2017; Xu et al., 2024) or JT-VAE (Jin et al., 2018).
-Fine-tuning
-Given an objective function $f\colon\mathcal{M}^{\prime}\to\mathbb{R}\cup\{-\infty\}$, we fine-tune the pretrained policy network using the learning algorithm described in Section 2.2.
-Unless stated otherwise, we initialize the molecule $m_{0}$ as a single carbon atom.
-We set $s=500$ as the number of top molecules to retain throughout training.
-In each epoch, we sample between $q_{\text{min}}=10\cdot 1024$ and maximum of $q_{\text{max}}=50\cdot 1024$ unique action sequences.
-At the end of each epoch, we train the network on 20 batches of size 64, sampled uniformly from the top 500 molecules.
-We intentionally keep the number of training batches per epoch relatively low for the network’s size to prevent premature overfitting in all considered cases.
-However, we observe that in most runs, training on more batches does not adversely affect performance and can actually increase the speed of convergence.
-3.4 Benchmark methods
-We compare our proposed approach with four diverse molecular design methods: Graph GA (Jensen, 2019) (a genetic algorithm (GA) applied to molecular graphs), REINVENT-Transformer (Xu et al., 2024) (a transformer model that generates SMILES strings), Junction Tree VAE (Jin et al., 2018) combined with a GA for optimization in latent molecular space, and STONED (Nigam et al., 2021) (which performs string mutations on the SELFIES (Krenn et al., 2020) representation).
-For benchmarking, we use the source code provided by Gao et al.
+Alphabet and size constraints
+We focus our solvent design task on organic H-C-N-O chemistry, restricting the available alphabet to $\Sigma=(\mathrm{C},\mathrm{N},\mathrm{O})$ without ionization and chirality.
+This choice emphasizes the primary building blocks and reduces the likelihood of designing unstable and unsustainable solvents by omitting elements more likely to contribute to these issues.
+Furthermore, to prevent potential exploitation of the surrogate by generating excessively large molecules that exceed the typical size of solvents, we constrain the design to molecules with no more than 25 atoms.
+Benchmark methods
+In addition to Graph GA and REINVENT-Transformer, we benchmark GraphXForm for solvent design against STONED (Nigam et al., 2021) and the Junction Tree Variational Autoencoder (JT-VAE) (Jin et al., 2018).
+For these methods, we use the source code provided by Gao et al.
 (2022).
-Although these methods employ different paradigms, we apply the same objective function across all methods to guide the design process.
-Graph GA and STONED
-Graph GA (Jensen, 2019) uses a GA that directly operates on the molecular graph by mutating atoms and fragments, using crossover rules derived from graph matching.
-It is a strong non-learning method that has been shown to outperform several SMILES-based learning approaches (Jensen, 2019; Gao et al., 2022).
-In contrast, STONED (Nigam et al., 2021) employs a GA that operates on the string level and manipulates tokens within the SELFIES molecular representation (Krenn et al., 2020).
-REINVENT-Transformer
-REINVENT-Transformer (Xu et al., 2024) is a recent state-of-the-art method for designing molecules by synthesizing SMILES strings in an autoregressive way.
-The approach involves a transformer network that is pretrained in a self-supervised manner on known molecules and then fine-tuned with reinforcement learning using a variant of the REINFORCE algorithm (Williams, 1992).
-REINVENT-Transformer is similar to GraphXForm in that both rely on pretrained transformers and generate molecules autoregressively.
-However, REINVENT-Transformer constructs molecules in a language-like manner by building a string token by token from an allowed vocabulary, whereas GraphXForm operates directly on the molecular graph by adding atoms and bonds.
-This direct manipulation of the molecular graph allows GraphXForm to guarantee the design of feasible molecules by construction, whereas REINVENT-Transformer must learn the SMILES grammar to avoid invalid structures.
-Additionally, our approach can start from arbitrary structures, unlike the left-to-right string construction in REINVENT-Transformer, making it straightforward to encode specific constraints to limit the search space.
-For a fair comparison, we pretrain the policy for REINVENT-Transformer using the same dataset as GraphXForm.
-Junction Tree VAE
-We further use variational autoencoders (VAEs) that have been widely used as a generative model in ML-guided molecular design, cf.
-overviews in Refs.
-(Sanchez-Lengeling & Aspuru-Guzik, 2018; Bilodeau et al., 2022; Anstine & Isayev, 2023).
-VAEs use an encoder-decoder structure: The encoder maps molecules to a continuous vector representation, referred to as latent space, from which the decoder reconstructs these molecules.
-The molecular latent space can facilitate exploration of the molecular space.
+STONED is a simple yet efficient algorithm that employs a GA operating at the string level, manipulating tokens within the SELFIES molecular representation (Krenn et al., 2020).
+JT-VAE adopts a VAE, a widely used generative model in ML-guided molecular design (Sanchez-Lengeling & Aspuru-Guzik, 2018; Bilodeau et al., 2022; Anstine & Isayev, 2023).
+VAEs use an encoder-decoder structure, where the encoder maps molecules into a continuous latent space, and the decoder reconstructs them from this representation.
+Particularly when the objective function is derived from a trained network, the molecular latent space can facilitate exploration of the molecular space.
 That is, different optimization strategies can be employed to discover points in the latent space that correspond to promising novel molecules.
 These strategies include random sampling, Bayesian optimization, and GAs (Sanchez-Lengeling & Aspuru-Guzik, 2018; Rittig et al., 2023d; Anstine & Isayev, 2023).
-We herein use the Junction Tree VAE (JT-VAE)  (Jin et al., 2018) in combination with GAs.
-The JT-VAE operates on molecular graphs and their non-cyclic abstractions (junction trees), and has shown a high rate of decoding vectors from its latent space to chemically valid molecules.
-For the solvent design task, we only consider molecules that conform to the alphabet $\Sigma$.
-We therefore train JT-VAE on a subset of the QM9 data set, which consists of about 128,000 molecules containing at most nine heavy atoms.
-We use GAs to optimize within the JT-VAE latent space because they enable exploring a large number of latent vectors, hence molecules, due to the linear scaling with the number of sampling points.
-3.5 Design results
-The goal for all methods is to find suitable solvents with respect to the two objectives from Equations 7 and 9 for the two example problems, using the general setup outlined in Section 3.2.
-Table 1: Performance of different molecular design methods for the two example tasks.
-Each method is run across three different random seeds, with a time budget of 8 hours.
+In our work, we use JT-VAE in combination with GAs.
+JT-VAE operates on molecular graphs and their non-cyclic abstractions (junction trees), and it has demonstrated a high rate of decoding latent vectors into chemically valid molecules.
+Since we consider only molecules that conform to the alphabet $\Sigma$, we train JT-VAE on a subset of the QM9 dataset (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) consisting of approximately 128,000 molecules with at most nine heavy atoms.
+3.3.1 Unconstrained results
+The goal for all methods is to find suitable solvents with respect to the two objectives from Equations 9 and 11 for the two example problems, using the general setup outlined in Section 3.3.
+In addition to the molecules shown in this section, a list of the top 20 molecules from the best runs can be found in Supplementary Figures 1-14.
+Table 2: Performance of different molecular design methods for the two example solvent design tasks.
+Each method is run across three different random seeds, with a maximum time budget of 8 hours.
 We report the objective function evaluation of the best molecule found over all runs (‘max best’), best averaged over all the three runs $\pm$ standard deviation (‘mean best’), the average of the top 20 molecules of the best run (‘max top 20’), and the mean of the top 20 over all three runs $\pm$ standard deviation (‘mean top 20’).
@@
 REINVENT-Transformer (Xu et al., 2024) 8.87 8.32 $\pm$ 0.52 8.66 8.09 $\pm$ 0.45 7.41 6.52 $\pm$ 1.17 7.22 6.42 $\pm$ 0.96
-GraphXForm (ours) 8.87 8.87 $\pm$ 0.00 8.67 8.67 $\pm$ 0.00 9.41 9.33 $\pm$ 0.12 9.23 9.14 $\pm$ 0.11
+GraphXForm (ours) 8.87 8.87 $\pm$ 0.00 8.60 8.58 $\pm$ 0.04 8.65 8.65 $\pm$ 0.00 8.41 8.39 $\pm$ 0.01
 Screening list
 To contextualize the results of all methods, we selected all 6,098 compounds from the COSMObase 2020 database that conform to the alphabet $\Sigma$.
-We calculated the objective functions for those compounds that also meet the miscibility gap requirements, cf.
+We calculated the objective functions for those compounds that also meet the miscibility gap requirements, see Eq.
 (8).
@@
 Additionally, Figures 3 and 4 display the structural formulas of the top three molecules identified by each method.
-Figure 3: IBA task: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
-Figure 4: TMB/DMBA task: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
+We note that, in these results, we do not yet impose any structural constraints and focus solely on optimization.
+Figure 3: IBA task, unconstrained: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
+Figure 4: TMB/DMBA task, unconstrained: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
 In the IBA task, GraphXForm, REINVENT-Transformer, and Graph GA all identified the same best molecule, which has an objective value of 8.87.
 However, GraphXForm consistently found this molecule in every run, highlighting its stability.
-For example, in contrast, while Graph GA achieved the same average value for the top 20 molecules, its mean best objective was only 7.13.
-For the more challenging TMB/DMBA task, GraphXForm significantly outperforms all others across every metric.
-Notably, Graph GA performs better overall than the recent REINVENT-Transformer, likely due to REINVENT-Transformer’s sensitivity to initialization during RL fine-tuning.
-This is also reflected in its relatively high standard deviation compared to the other methods.
+For example, in contrast, while Graph GA has a slightly higher average value for the best 20 molecules found, its mean best objective over all runs is only 7.13.
+For the more challenging TMB/DMBA task, GraphXForm outperforms all other methods across every metric.
+Notably, Graph GA outperforms REINVENT-Transformer overall, likely due to the latter’s sensitivity to initialization during RL fine-tuning.
+This factor is also reflected in its relatively high standard deviation.
+We also observe that the designs produced by Graph GA and GraphXForm exhibit substantial structural similarity, although GraphXForm makes additional refinements that lead to improved scores.
 Interestingly, the JT-VAE and STONED methods identified molecules with significantly lower objective values for the TMB/DMBA task when compared to other methods, with maximum values of approximately 2.16 and 2.39, respectively.
 However, their results for the IBA task (6.85 and 7.53) were closer to those of the other methods.
-We attribute this discrepancy to the nature of the tasks: The TMB/DMBA task is inherently more challenging and requires larger, more complex molecules, while the best-performing molecules in the IBA task were smaller.
+We attribute this discrepancy to the nature of the tasks: The TMB/DMBA task is inherently more challenging and requires larger molecules with more complex branching, while the best-performing molecules in the IBA task were smaller.
 As JT-VAE was trained only on molecules with up to nine heavy atoms, its ability to explore larger molecules seems limited.
 Similarly, STONED struggled to effectively explore larger molecular structures.
-In summary, GraphXForm demonstrates highly promising results in both solvent design tasks, outperforming state-of-the-art molecular design methods in terms of identifying the best candidates and ensuring robustness in the design process.
+In summary, GraphXForm demonstrates highly promising results in both solvent design tasks, outperforming its comparison partners in terms of identifying the best candidates and ensuring robustness in the design process.
 In the following sections, we further explore the flexibility of GraphXForm by imposing structural constraints on the designed molecules and starting the design process from initial molecular structures.
-3.6 Including structural constraints in molecular design
+3.3.2 Including structural constraints in molecular design
 In addition to optimizing an objective, such as a physical property, it is often desirable to enforce specific structural characteristics in the generated molecular candidates.
 These structural constraints can include limitations on ring size or specific bonding patterns between atom types, which can improve chemical feasibility, such as enhancing synthesizability and stability or reducing toxic moieties.
-For instance, in the TMB/DMBA task, although the top molecules generated by GraphXForm exhibit high objective values, they possess several characteristics that pose challenges for solvent design.
-Notably, the presence of single nitrogen-nitrogen bonds often leads to instability (hyd, 2024), resulting in reactive compounds that are unsuitable for liquid-liquid extraction.
-Additionally, the existence of a nine-member ring could pose significant challenges to the solvent’s synthesizability (Steinborn et al., 2023).
-As noted in Section 2.1, string-based molecular design methods cannot directly enforce such structural constraints.
-Typically, this is managed by assigning a very low objective value to molecules that violate these constraints.
-In contrast, by operating directly on the molecular graph, GraphXForm allows us to restrict the search space for many constraints by simply masking actions that would lead to constraint-violating molecules.
-To examine how structural constraints affect the performance of GraphXForm in finding molecules with high objective values, we impose the following constraints: We limit the ring size to a maximum of six atoms and disallow any bonds between two oxygen atoms as well as single bonds between two nitrogen atoms.
-We note that these constraints of course do not cover all possibilities for maintaining synthetic accessibility and chemical stability; instead, they serve as examples to illustrate the capabilities of GraphXForm.
-For the IBA task, the best molecule identified in Table 1 with objective value of 8.87 contains multiple nitrogen-nitrogen single bonds.
-We therefore rerun the design task with the additional structural constraints.
-The top three molecules designed by GraphXForm in these runs are shown in the top row of Figure 5.
-With constraints, we achieve a ‘mean best’ value of $8.23\pm 0.00$ and a ‘mean top 20’ value $7.98\pm 0.02$, which are only slightly lower than the values obtained without structural constraints.
-We also repeated these constrained design runs with REINVENT-Transformer, given its conceptual similarity to GraphXForm.
-However, REINVENT-Transformer only reached a ‘mean best’ value of $7.03\pm 1.14$, with its overall best molecule having an objective value of 7.87.
-The top three molecules identified by GraphXForm with the structural constraints for the TMB/DMBA task are shown in the bottom row of Figure 5.
-Under these constraints, GraphXForm achieves a ‘mean best’ value of $8.65\pm 0.00$ and a ‘mean top 20’ value of $8.42\pm 0.03$.
-Notably, these results remain superior to those of the other methods that do not impose structural constraints listed in Table 1.
-This further demonstrates GraphXForm’s flexibility and its ability to consistently design promising molecules.
-Figure 5: Top three molecules designed by GraphXForm where: ring size is limited to a maximum of six atoms, any bond between oxygen atoms is disallowed, and single bonds between nitrogen atoms are disallowed.
-3.7 Starting molecule design from initial structures
-Figure 6: Best molecule designed by GraphXForm when adding a maximum of 3, 5, and 7 atoms to a specific initial structure.
-We note that increasing the order of a bond within the initial structure is allowed.
-For TMB/DMBA, we also enforce that the designed molecule stays an alcohol.
-One advantage of GraphXForm, which is not feasible when constructing SMILES strings via next-token prediction due to the linear construction, is the ability to start the molecule design process from a predefined structure.
-This feature is particularly useful when a known candidate, potentially in practical use, possesses desirable properties but could benefit from slight modifications to improve the objective function.
-To demonstrate this, we consider the IBA task and begin with the best molecule from COSMObase 2020, which has an objective value of 5.57 (see Figure 4).
-We conduct three individual experiment runs, allowing GraphXForm to add up to 3, 5, and 7 atoms to the initial structure (and alter bond orders as needed).
-We also maintain the structural constraints outlined in Section 3.6.
-The original molecule, along with the best molecule from each of the three runs, is shown in the first row of Figure 6.
-In all cases, the designed molecules exhibit improvements in the objective value.
+For example, in the IBA task, although the top molecules generated by GraphXForm achieve high objective values, they exhibit features that pose challenges for solvent design.
+Notably, the presence of single nitrogen–nitrogen bonds and single nitrogen–carbon–nitrogen bonds may lead to instability in water (hyd, 2024), resulting in reactive compounds unsuitable for liquid–liquid extraction.
+Similarly, our top result for the TMB/DMBA task includes a carbon atom that is single-bonded to a nitrogen, a hydroxyl group, and a single hydrogen atom, which also raises concerns about instability.
+A straightforward approach to enforcing such constraints is to assign an objective value of $\infty$- ∞ to any designed molecules that violate them, effectively discarding those candidates.
+In contrast, because GraphXForm operates directly on the molecular graph – transitioning from one molecule to the next without waiting for the completion of a string – we can flexibly restrict the search space by simply masking actions that would lead to constraint violations.
+To examine how structural constraints affect the performance of GraphXForm in finding molecules with high objective values, we impose the following constraints: rings are limited to five or six atoms, and the following bonding patterns are disallowed: single bonds between two nitrogen atoms; single bonds between two oxygen atoms; a carbon atom single-bonded to two nitrogen atoms, unless the carbon is also double-bonded to an oxygen atom, forming a urea functional group; and a carbon atom single-bonded to a nitrogen atom, an oxygen atom, a hydrogen atom, and one other non-hydrogen functional group.
+We acknowledge that these constraints do not cover all aspects of synthetic accessibility, chemical stability, safety, or environmental impact.
+Rather, they serve as illustrative examples to demonstrate the capabilities of GraphXForm in incrementally limiting the designs to promising candidate structures that can be further evaluated by experts.
+For the IBA task, the top three molecules designed by GraphXForm under these constraints are shown in the top row of Figure 5.
+With constraints in place, we achieve a ‘mean best’ value of $7.28\pm 0.00$ and a ‘mean top 20’ value $6.85\pm 0.02$, which are only slightly lower than the values obtained without structural constraints.
+We also conducted constrained design runs with REINVENT-Transformer as a string-based counterpart; however, it reached only a ‘mean best’ value of $6.36\pm 0.39$, with its overall best molecule scoring $6.92$.
+The bottom row of Figure 5 displays the top three molecules identified by GraphXForm under the structural constraints for the TMB/DMBA task.
+Under these constraints, GraphXForm achieves a ‘mean best’ value of $8.54\pm 0.00$ and a ‘mean top 20’ value of $8.31\pm 0.06$.
+Notably, these results remain superior to those of the other methods without structural constraints, as reported in Table 2.
+These findings further demonstrate GraphXForm’s flexibility and its ability to consistently design promising molecules.
+Figure 5: Top three molecules designed by GraphXForm under structural constraints on specific bonding patterns and ring sizes.
+3.3.3 Starting molecule design from initial structures
+Figure 6: Best molecule designed by GraphXForm when augmenting a specific initial structure by adding up to 3, 5, or 7 atoms.
+Bonds can be added without restriction.
+For the TMB/DMBA task, the designed molecule is also required to remain an alcohol.
+One advantage of our molecular graph approach, which is not achievable out of the box with vanilla SMILES-based next-token prediction due to its linear construction, is the direct ability to initiate the design process from a predefined structure.
+This feature is especially useful when a known candidate, which already possesses desirable properties, could benefit from targeted modifications to further improve the objective function.
+To demonstrate this, we consider the IBA task and start with the best molecule from COSMObase 2020, which has an objective value of 5.57 (see Figure 4).
+We then conduct three separate experimental runs, allowing GraphXForm to add up to 3, 5, and 7 atoms to the initial structure while adding bonds as needed, all under the structural constraints outlined in Section 3.3.2.
+The original molecule and the best molecule from each of the three runs are shown in the top row of Figure 6.
+In every case, the same modified molecule is produced (with three additional atoms), resulting in an improved objective value.
 We repeat a similar experiment for the TMB/DMBA task.
-During the list screening, we observed that the top three structures were all long-chain alcohols.
-However, long-chain alcohols tend to have melting points above room temperature, making them unsuitable as solvents for the processes under consideration.
-The melting point can be lowered by adding branches and increasing bond orders.
-Thus, we initiate three runs starting from the best molecule identified in the list screening, allowing the addition of up to 3, 5, or 7 atoms, respectively.
-Additionally, we constrain the design process to ensure the resulting molecule remains an alcohol by preventing any actions from being taken on the hydroxy group (one oxygen bonded to one hydrogen).
-The results are shown in the second row of Figure 6.
-While no better molecule is found when adding only up to 3 atoms, a notable improvement occurs when allowing to add 5 or 7 atoms.
-This approach provides a powerful tool for situations where it is preferable not to start the design process from scratch but to make targeted modifications to existing molecules.
-We note that allowing the agent to remove atoms or bonds would not provide any benefit when starting the design from scratch, i.e., from a single atom.
-In fact, it would introduce additional redundancies into the search space.
-However, in the discussed scenarios where the design begins with an existing molecule, enabling the agent to remove atoms or bonds could in fact offer an advantage, as it allows for greater flexibility and deviations from the initial structure.
-We leave the exploration of this possibility for future work.
+During list screening, we observed that the top three structures were all long-chain alcohols.
+However, long-chain alcohols tend to have melting points above room temperature, making them unsuitable as solvents for the intended processes.
+The melting point can be lowered by, e.g., adding branches.
+To address this, we again initiated three runs starting from the best molecule identified in the screening.
+These runs allowed the addition of up to 3, 5, and 7 atoms, thereby simulating branching of the long-chain alcohol.
+Additionally, we constrained the design process so that the resulting molecule remains an alcohol by preventing any modifications to the hydroxy group (an oxygen atom bonded to one hydrogen atom).
+The results, shown in the second row of Figure 6, indicate that while no improvement in the objective function is observed when adding only 3 atoms, a notable enhancement occurs when 5 or 7 atoms are added.
+This approach provides a powerful tool for cases where it is preferable to build upon an existing molecule rather than starting from scratch.
+We note that enabling the agent to remove atoms or bonds would not be beneficial when designing from a single atom, as it would introduce unnecessary redundancies into the search space.
+However, in scenarios where the design process begins with an existing molecule, allowing removal could offer additional flexibility and enable greater deviations from the initial structure.
+Extending the action space to include atom and bond removal is straightforward within our framework; however, we leave the exploration of this possibility for future work.
 Conclusion
@@
 By doing so, we addressed challenges faced by string-based methods, such as chemical validity or accomodating structural constraints.
-We introduced a technique derived from self-improvement learning to stably fine-tune a deep graph-transformer model with eight layers.
-On two solvent design tasks, we showed that GraphXForm can outperform state-of-the-art molecule design techniques.
+We introduced a technique derived from self-improvement learning to fine-tune a deep graph-transformer model.
+Based on the established drug design GuacaMol benchmark (Brown et al., 2019) and two solvent design tasks, we showed that GraphXForm can outperform state-of-the-art molecule design techniques.
 Additionally, our approach can flexibly adhere to specified structural constraints, such as bond types and functional groups, and can adaptively start the design process from existing molecular structures.
 Looking ahead, several promising avenues for future development exist.
-First, we could expand GraphXForm to include more atoms and ionization states, as well as pretrain the network on much larger databases.
-We note that these features can already be integrated within our current framework; however, we restricted the set of possible atoms to ensure comparibility between our and other methods.
-We could further evaluate GraphXForm on additional solvent design tasks, as well as more extensively researched drug design tasks.
-We also aim to incorporate more constraints to ensure the suitability of solvents for liquid-liquid separation processes, recognizing that numerous factors – such as boiling and melting points – are critical to a solvent’s effectiveness.
+First, we plan to expand the atom alphabet of GraphXForm and pretrain the network on significantly larger databases.
+Although these enhancements can be integrated into our current framework, we intentionally limited the atom set in this study to ensure comparability with other methods.
+Second, we aim to extend the action space by allowing the agent to remove bonds and atoms.
+While not necessary for our present investigation – and admittedly introducing further action symmetries – this addition would offer the model more possibilities when modifying starting molecules.
+Additionally, future work could involve evaluating GraphXForm on a broader range of molecular design tasks and incorporating more design considerations.
+For instance, recent trends in ML-based drug development emphasize increased sample efficiency (Gao et al., 2022; Guo & Schwaller, 2024a, c; Rittig et al., ) and the integration of synthesizability models (Guo & Schwaller, 2024b; Gao et al., 2024).
+Although sample efficiency was not the primary focus of our current study, our method inherently maintains a kind of experience replay buffer, suggesting that techniques targeting replay buffers (Guo & Schwaller, 2024a, c) could be readily applied to further enhance GraphXForm.
+Furthermore, tweaking TASAR parameters (e.g., using a shallower beam width $\beta$ or a higher step size $\sigma$) and allowing more training batches in each fine-tuning epoch can help to make the method more sample-efficient.
+We also aim to incorporate more constraints to ensure the suitability of the designed molecules for specific applications.
+In the case of solvents for liquid-liquid separation processes, this includes recognizing that numerous factors – such as boiling and melting points – are critical to a solvent’s effectiveness.
 This, however, will depend on the availability of reliable property predictors.
@@
 The authors gratefully acknowledge the Competence Center for Digital Agriculture (KoDA) at the University of Applied Sciences Weihenstephan-Triesdorf for providing additional computational resources.
+Furthermore, the authors thank Herbert Riepl for suggesting structural constraints.
+Appendix A Atom alphabet
+We list the atom alphabet for the different design tasks as a Python dictionary.
+We also provide the valence of an atom, as well as its formal_charge and chiral_tag – for atom specification in RDKit – if needed.
+A.1 Drug design tasks
+        "C": {"atomic_number": 6, "valence": 4},
+        "C-": {"atomic_number": 6, "valence": 3, "formal_charge": -1},
+        "C+": {"atomic_number": 6, "valence": 5, "formal_charge": 1},
+        "C@": {"atomic_number": 6, "valence": 4, "chiral_tag": CHI_TETRAHEDRAL_CW},
+        "C@@": {"atomic_number": 6, "valence": 4, "chiral_tag": CHI_TETRAHEDRAL_CCW},
+        "N": {"atomic_number": 7, "valence": 3},
+        "N-": {"atomic_number": 7, "valence": 2, "formal_charge": -1},
+        "N+": {"atomic_number": 7, "valence": 4, "formal_charge": 1},
+        "O": {"atomic_number": 8, "valence": 2},
+        "O-": {"atomic_number": 8, "valence": 1, "formal_charge": -1},
+        "O+": {"atomic_number": 8, "valence": 3, "formal_charge": 1},
+        "F": {"atomic_number": 9, "valence": 1},
+        "P": {"atomic_number": 15, "valence": 7},
+        "P-": {"atomic_number": 15, "valence": 6, "formal_charge": -1},
+        "P+": {"atomic_number": 15, "valence": 8, "formal_charge": 1},
+        "S": {"atomic_number": 16, "valence": 6},
+        "S-": {"atomic_number": 16, "valence": 5, "formal_charge": -1},
+        "S+": {"atomic_number": 16, "valence": 7, "formal_charge": 1},
+        "S@": {"atomic_number": 16, "valence": 6, "chiral_tag": CHI_TETRAHEDRAL_CW},
+        "S@@": {"atomic_number": 16, "valence": 6, "chiral_tag": CHI_TETRAHEDRAL_CCW},
+        "Cl": {"atomic_number": 17, "valence": 1},
+        "Br": {"atomic_number": 35, "valence": 1},
+        "I": {"atomic_number": 53, "valence": 1}
+A.2 Solvent design tasks
+        "C": {"atomic_number": 6, "valence": 4},
+        "N": {"atomic_number": 7, "valence": 3},
+        "O": {"atomic_number": 8, "valence": 2}
+Appendix B Drug design results
+Here, we present the complete results for the 20 goal-directed tasks of the GuacaMol benchmark.
+To put our results into even further perspective, we included additional baseline methods not mentioned in the main paper.
+Table 3: GuacaMol goal-directed benchmarks
+Benchmark SMILES GA‡ [1] SMILES LSTM‡ [2] Graph GA‡ [3] Reinvent¶ [4] GEGL¶ [5] CReM [6] MolRL-MGPT [7] GraphXForm (ours)
+Celecoxib rediscovery 0.732 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Troglitazone rediscovery 0.515 1.000 1.000 1.000 0.552 1.000 1.000 1.000
+Thiothixene rediscovery 0.598 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Aripiprazole similarity 0.834 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Albuterol similarity 0.907 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Mestranol similarity 0.790 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+C11H24 * 0.829 0.993 0.971 0.999 1.000 0.966 1.000 1.000
+C9H10N2O2PF2Cl † 0.889 0.879 0.982 0.877 1.000 0.940 0.939 1.000
+Median molecules 1 0.334 0.438 0.406 0.434 0.455 0.371 0.449 0.459
+Median molecules 2 0.380 0.422 0.432 0.395 0.437 0.434 0.422 0.403
+Osimertinib MPO 0.886 0.907 0.953 0.889 1.000 0.995 0.977 0.972
+Fexofenadine MPO 0.931 0.959 0.998 1.000 1.000 1.000 1.000 0.999
+Ranolazine MPO 0.881 0.855 0.920 0.895 0.933 0.969 0.939 0.944
+Perindopril MPO 0.661 0.808 0.792 0.764 0.833 0.815 0.810 0.835
+Amlodipine MPO 0.722 0.894 0.894 0.888 0.905 0.902 0.906 0.923
+Sitagliptin MPO 0.689 0.545 0.891 0.539 0.749 0.763 0.823 0.965
+Zaleplon MPO 0.413 0.669 0.754 0.590 0.763 0.770 0.790 0.727
+Valsartan SMARTS 0.552 0.978 0.990 0.095 1.000 0.994 0.997 1.000
+Deco hop 0.970 0.996 1.000 0.994 1.000 1.000 1.000 1.000
+Scaffold hop 0.885 0.998 1.000 0.990 1.000 1.000 1.000 1.000
+Total score 14.398 17.341 17.983 16.349 17.627 17.919 18.052 18.227
+* Top-159, no stereochemistry. † Top-250, no stereochemistry. ‡ Results as reported in [8]. ¶ Results as reported in [7].
+[1] Naruki Yoshikawa, Kei Terayama, Masato Sumita, Teruki Homma, Kenta Oono, and Koji Tsuda.
+Population-based de novo molecule generation, using grammatical evolution.
+Chemistry Letters, 47(11):1431–1434, 2018.
+[2] Marwin HS Segler, Thierry Kogej, Christian Tyrchan, and Mark P Waller.
+Generating focused molecule libraries for drug discovery with recurrent neural networks.
+ACS central science, 4(1):120–131, 2018.
+[3] Jan H Jensen. A graph-based genetic algorithm and generative model/monte carlo tree search for the exploration of chemical space. Chemical science, 10(12):3567–3572, 2019.
+[4] Marcus Olivecrona, Thomas Blaschke, Ola Engkvist, and Hongming Chen. Molecular de-novo design through deep reinforcement learning. Journal of Cheminformatics, 9, 2017.
+[5] Sungsoo Ahn, Junsu Kim, Hankook Lee, and Jinwoo Shin. Guiding deep molecular optimization with genetic exploration. Advances in Neural Information Processing Systems, 33, 12008–12021, 2020.
+[6] Pavel Polishchuk. CReM: chemically reasonable mutations framework for structure generation. Journal of Cheminformatics, 12, 2020.
+[7] Xiuyuan Hu, Guoqing Liu, Yang Zhao, and Hao Zhang.
+De novo drug design using reinforcement learning with multiple gpt agents.
+Advances in Neural Information Processing Systems, 36, 7405-7418, 2023.
+[8] Brown, N.; Fiscato, M.; Segler, M. H. S.; Vaucher, A. C. GuacaMol: Benchmarking Models for de Novo Molecular Design. Journal of Chemical Information and Modeling, 59(3):1096–1108, 2019.
+Appendix C Solvent design results
+C.1 GraphXForm
+Figure 7: Top 20 molecules: GraphXForm, IBA task
+Figure 8: Top 20 molecules: GraphXForm, TMB/DMBA task
+C.2 List screen
+Figure 9: Top 20 molecules: List screen, IBA task
+Figure 10: Top 20 molecules: List screen, TMB/DMBA task
+C.3 JT-VAE
+Figure 11: Top 20 molecules: JT-VAE, IBA task
+Figure 12: Top 20 molecules: JT-VAE, TMB/DMBA task
+C.4 STONED
+Figure 13: Top 20 molecules: STONED, IBA task
+Figure 14: Top 20 molecules: STONED, TMB/DMBA task
+C.5 Graph GA
+Figure 15: Top 20 molecules: Graph GA, IBA task
+Figure 16: Top 20 molecules: Graph GA, TMB/DMBA task
+C.6 REINVENT-Transformer
+Figure 17: Top 20 molecules: REINVENT-Transformer, IBA task
+Figure 18: Top 20 molecules: REINVENT-Transformer, TMB/DMBA task
+C.7 GraphXForm: Structural constraints
+Figure 19: Top 20 molecules: GraphXForm, IBA task with structural constraints
+Figure 20: Top 20 molecules: GraphXForm, TMB/DMBA task with structural constraints